### v2

# On Weak Compactness and Uniform Regularity in Lipschitz free Spaces

Abstract.
We analyze the properties of weakly compact sets in Lipschitz free spaces. Prior research has established that, for a complete metric space $M$, weakly precompact sets in the Lipschitz free space $\mathcal{F}(M)$ are tight. In this paper, we prove that these sets actually exhibit a stronger property, which we call uniform regularity. However, this condition alone is not sufficient to characterize weakly compact sets, except in the case of scattered metric spaces. On the other hand, if $T$ is an $\mathbb{R}$-tree, we leverage Godard’s isometry between $\mathcal{F}(T)$ and $L^{1}(\lambda_{T})$ to obtain an intrinsic characterization of weakly compact sets in $\mathcal{F}(T)$. This approach allows us to identify conditions that may describe weak compactness across a wider range of spaces. In particular, we provide a characterization of norm-compactness in terms of sums of “large molecules”, while we show that sums of “small molecules” contain an $\ell_{1}$-basis.
Key words and phrases:
Lipschitz free space, weakly compact set, V*-set, tightness, uniform regularity
2020 Mathematics Subject Classification:
Primary 46B50, 46B20
Introduction
The weak topology is of the utmost importance when studying the geometry of Banach spaces. In particular, a good knowledge of weakly compact subsets is crucial. Indeed, many Banach space properties are intimately related to weak compactness, which in turn provide very helpful tools to understand, analyze, and classify Banach spaces. In this paper, we study weakly compact sets in the so-called Lipschitz free spaces. In a nutshell, for a metric space $(M,d)$, the Lipschitz free space $\mathcal{F}(M)$ is the free vector space generated by $M$, equipped with a particular norm which makes it a Banach space. The norm is designed in such a way that $M$ is isometric to a linearly dense subset $\delta(M)$ of $\mathcal{F}(M)$, and moreover, Lipschitz maps from $\delta(M)$ into any Banach space $X$ uniquely extend to bounded linear operators from $\mathcal{F}(M)$ into $X$. See Section 1.2 for more details.
Weak convergence and weakly compact sets in general Lipschitz free spaces are relatively poorly understood. Nevertheless, a successful first step was made in [5] by isolating a powerful necessary condition, called tightness. Recall that a set of measures $S$ on $M$, equipped with the Borel $\sigma$-algebra, is tight if, for any $\varepsilon>0$, there is a compact subset $K\subset M$ such that $|\mu(M\setminus K)|<\varepsilon$ for all measures $\mu\in S$ ($|\mu|$ being the total variation of $\mu$). By analogy to this notion, a bounded subset $W$ in a Lipschitz free space $\mathcal{F}(M)$ is called tight if, for any $\varepsilon>0$, there is a compact subset $K\subset M$ such that:
$W\subset\mathcal{F}(K)+\varepsilon B_{\mathcal{F}(M)},$
where $B_{\mathcal{F}(M)}$ is the unit ball of $\mathcal{F}(M)$. Roughly speaking, this says that every element of $W$ can be approximated arbitrarily close by members of a Lipschitz free space over a compact subset $K$ of $M$. It is proved in [5] that “Every weakly compact subset of a free space is tight”. This of course bears a resemblance to Prokhorov’s theorem, which asserts that a collection $S$ of probability measures is tight if and only if the closure of $S$ is sequentially compact in the space of probability measures equipped with the topology of weak convergence of measures. Unfortunately, tightness in Lipschitz free spaces is far from being sufficient for weak compactness. For instance, for any infinite compact metric space $M$, $B_{\mathcal{F}(M)}$ is tight while obviously not weakly compact (infinite-dimensional free spaces are never reflexive). Nevertheless, the discovery of this condition had an immediate effect on the general understanding of linear properties such as the Schur property (SP), weak sequential completeness (wsc), the Radon-Nikodým property (RNP), and the approximation property (AP), to mention a few. See e.g. [4, 5, 9, 15].
The first half of this paper deals with a strengthening of the tightness condition, called uniform regularity. Let us denote $\mathcal{M}(K)$ the space of all Radon (i.e. finite and regular) signed Borel measures on a compact Hausdorff space $K$. Recall that a subset $S\subset\mathcal{M}(K)$ is said to be uniformly regular if given any open set $U\subset K$ and $\varepsilon>0$, there is a compact set $H\subset U$ such that $|\mu(U\setminus H)|<\varepsilon$ for all $\mu\in S$. A famous result due to Grothendieck (see e.g. [3, Theorem 5.3.2]) states that a bounded subset $S\subset\mathcal{M}(K)$ is relatively weakly compact if and only if it is uniformly regular. By analogy, we introduce the following definition:
Definition.
Let $M$ be a pointed metric space. We say that a bounded subset $W\subset\mathcal{F}(M)$ is uniformly regular if for any open set $U\subset M$ and $\varepsilon>0$, there is a compact set $K\subset U$ such that:
$W\subset\mathcal{F}(K\cup(M\setminus U))+\varepsilon B_{\mathcal{F}(M)}.$
It is rather easy to see that uniform regularity is strictly stronger than tightness (see Remark 2.2). Our first goal will be to prove the next result:
Theorem A.
Let $M$ be a complete pointed metric space. If $W\subset\mathcal{F}(M)$ is a V*-set, in particular if $W$ is relatively weakly compact, then it is uniformly regular.
See Section 1.3 for the definition of V*-sets. Unfortunately, uniform regularity alone is not sufficient to characterize weak compactness for a general $M$. However, it does so on every complete scattered metric space.
Theorem B.
Let $M$ be a complete metric space. The following assertions are equivalent:
  1. (1)
    $M$ is scattered;
  2. (2)
    Every uniformly regular set in $\mathcal{F}(M)$ is relatively compact;
  3. (3)
    Every uniformly regular set in $\mathcal{F}(M)$ is relatively weakly compact.
  4. (4)
    Every uniformly regular set in $\mathcal{F}(M)$ is a V*-set.
In particular, the above theorem shows that for $M=C$, the middle-thirds Cantor set, there exists a uniformly regular set that is not relatively weakly compact.This is notable because $\mathcal{F}(C)$ enjoys the Schur property (being isometric to $\ell_{1}$, see e.g. [16]) and so weak compactness and norm compactness are actually equivalent in this setting. It also yields the following corollary, which appears to be new.
Corollary C.
If $M$ is a complete scattered metric space, then $\mathcal{F}(M)$ has property (V*); that is, every V*-set in $\mathcal{F}(M)$ is relatively weakly compact.
In the second half of the paper we tackle the problem from a different angle. Recall that an $\mathbb{R}$-tree is an arc-connected metric space $(T,d)$ with the property that there is a unique arc connecting any pair of points $x\neq y\in T$ and it is moreover isometric to the real segment $0,d(x,y)]\subset\mathbb{R}$. In such a tree $T$, there is a canonical measure $\lambda_{T}$, called the length measure, which provides a natural way to measure the size of subsets, similar to Lebesgue measure on the real line. Godard proved in [16] that $\mathcal{F}(T)$ can be naturally identified with the space $L^{1}(\lambda_{T})$. Now one should keep in mind that relatively weakly compact sets in $L^{1}(\lambda_{T})$ are well understood: they are precisely the uniformly integrable subsets of $L^{1}(\lambda_{T})$. Therefore we will transfer the equi-integrability conditions through the isometry $L^{1}(\lambda_{T})\to\mathcal{F}(T)$ and describe the resulting properties in $\mathcal{F}(T)$. This will provide us some new insights on the properties of weakly compact sets. More precisely, this highlights the importance of representation of elements in $W$ as combinations of elementary molecules satisfying specific properties. We remind the reader that every element $\gamma$ in $\mathcal{F}(M)$ can be written as a series $\gamma=\sum_{i=1}^{\infty}a_{i}m_{x_{i}y_{i}}$ where $(a_{i})_{i}\in\ell_{1}$ and $m_{x_{i}y_{i}}=d(x_{i},y_{i})^{-1}(\delta(x_{i})-\delta(y_{i}))$ are molecules. If $\|\gamma\|=\sum_{i=1}^{\infty}|a_{i}|$, such a representation is then referred to as convex series of molecules.
Theorem D.
Let $T$ be an $\mathbb{R}$-tree. If $W\subset\mathcal{F}(T)$ then $W$ is relatively weakly compact if and only if it is bounded and satisfies these two properties:
  1. $(a)$
    $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every convex series of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$ with $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant\varepsilon$, $\forall I\subset\{1,\ldots,n\}$, $\sum_{I}d(x_{i},y_{i})<\delta$ implies that $\sum_{I}|a_{i}|<2\varepsilon$.
  2. $(b)$
    For any $\varepsilon>0$, there exists a finitely generated subtree $T^{\prime}:=\bigcup_{i=1}^{n}[0,z_{i}]\subset T$ such that $W\subset\mathcal{F}(T^{\prime})+\varepsilon B_{\mathcal{F}(T)}$.
Informally, property (a) says that the weight of molecules formed by pairs of points in close proximity should be controlled when approximating elements of a weakly compact set by convex series of molecules. Our next result shows that this is somehow necessary in general.
Theorem E.
Let $M$ be a complete pointed metric space. For every $n\in\mathbb{N}$, let $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\in S_{\mathcal{F}(M)}$ be such that $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ and $\sup_{n}\sum_{k}\lambda_{k}^{n}<\infty$. Then $(\gamma_{n})_{n}$ has a subsequence which is an (asymptotically isometric) $\ell_{1}$-basis.
Clearly, if a set $W$ contains a sequence $(\gamma_{n})$ as above, then $W$ is not relatively weakly compact.
1. Preliminaries
1.1. Notation
The letter $X$ is reserved for a real Banach space. As usual, $X^{*}$ denotes the topological dual of $X$, while $B_{X}$ and $S_{X}$ stand for the unit ball and the unit sphere of $X$, respectively. If $S\subset X$ then $\operatorname{conv}(S)$ is the convex hull of $S$. Ordinarily, $L^{1}=L^{1}([0,1])$ represents the Banach space of Lebesgue integrable functions from $[0,1]$ to $\mathbb{R}$. Lebesgue measure in $\mathbb{R}$ is denoted by $\lambda$. Similarly, $\ell_{1}=\ell_{1}(\mathbb{N})$ is the space of absolutely summable sequences in $\mathbb{R}$ and $c_{0}$ is the space of all real sequences which converge to 0. To indicate that $X$ is isometrically isomorphic to another Banach space $Y$, we use the notation $X\equiv Y$.
Throughout the paper, $(M,d)$ is a complete and pointed metric space, with distinguished point denoted by 0. If $p\in M$ and $r\geqslant 0$, then we will use the notation:
$B(p,r)$ $=\{x\in M\;|\;d(x,p)\leqslant r\}$
$B^{O}(p,r)$ $=\{x\in M\;|\;d(x,p)<r\}$
Moreover, for $S\subset M$ we write $S^{c}:=M\setminus S$ and
$[S]_{r}$ $=\left\{{x\in M:d(x,S)\leqslant r}\right\}$
$\widetilde{S}$ $=\{(x,y)\in S\times S\;:\;x\neq y\}.$
If $M,N$ are pointed metric spaces, with both base points denoted by 0, then $\operatorname{Lip}_{0}(M,N)$ stands for the set of all Lipschitz functions $f:M\to N$ such that $f(0)=0$. As usual, $\operatorname{Lip}(f)$ refers to the best Lipschitz constant of $f$. If $N=X$ is a real Banach space, then $\operatorname{Lip}_{0}(M,X)$ naturally becomes a Banach space when equipped with the Lipschitz norm:
$\forall f\in\operatorname{Lip}_{0}(M,X),\quad\|f\|_{L}:=\operatorname{Lip}(f)=\sup_{x\neq y\in M}\frac{\|f(x)-f(y)\|_{X}}{d(x,y)}.$
In the case $X=\mathbb{R}$, it is customary to write $\operatorname{Lip}_{0}(M):=\operatorname{Lip}_{0}(M,\mathbb{R})$.
1.2. Lipschitz free spaces
For $x\in M$, we let $\delta(x)\in\operatorname{Lip}_{0}(M)^{*}$ be the evaluation functional defined by $\langle\delta(x),f\rangle=f(x),\ \forall f\in\operatorname{Lip}_{0}(M)$. It is readily seen that $\delta:x\in M\mapsto\delta(x)\in\operatorname{Lip}_{0}(M)^{*}$ is an isometry. The Lipschitz free space over $M$ is then defined as the norm-closure of the space generated by such evaluation functionals:
$\mathcal{F}(M):=\overline{\mbox{span}}^{\|\cdot\|}\left\{\delta(x)\,:\,x\in M\right\}\subset\operatorname{Lip}_{0}(M)^{*}.$
It is well known that $\mathcal{F}(M)$ is actually a predual of $\operatorname{Lip}_{0}(M)$, that is $\operatorname{Lip}_{0}(M)\equiv\mathcal{F}(M)^{*}$. For an in-depth understanding of Lipschitz free spaces, interested readers are referred to [25] (where they are called Arens-Eells spaces).
Let us recall some important features of Lipschitz free spaces. Firstly, note that, on bounded subsets of $\operatorname{Lip}_{0}(M)$, the weak topology induced by $\mathcal{F}(M)$ coincides with the pointwise topology. If $N\subset M$ is closed, then $\mathcal{F}(N\cup\left\{{0}\right\})\equiv\mathcal{F}_{M}(N)$ where:
$\mathcal{F}_{M}(N):=\overline{\text{span}}\{\delta_{M}(x)\;|\;x\in N\}\subset\mathcal{F}(M).$
Under this identification, the support of $\gamma\in\mathcal{F}(M)$ is the smallest closed subset $N\subset M$ such that $\gamma\in\mathcal{F}_{M}(N\cup\left\{{0}\right\})$. It is denoted by $\text{supp}(\gamma)$. The set of finitely supported elements of $\mathcal{F}(M)$ is precisely $\mathrm{span}\,\delta(M)$, so it is dense in $\mathcal{F}(M)$. For further details, we refer to [7, 8].
Next, let $f:M\to N$ be a Lipschitz map preserving the basepoint. By the universal linearization property of Lipschitz-free spaces, there exists a unique bounded linear operator
$\widehat{f}:\mathcal{F}(M)\to\mathcal{F}(N)$
satisfying $\widehat{f}\circ\delta_{M}=\delta_{N}\circ f$. Moreover, $\|\widehat{f}\|=\operatorname{Lip}(f)$.
If $x\neq y\in M$, then the elementary molecule $m_{xy}\in S_{\mathcal{F}(M)}$ is defined by:
$m_{xy}:=\frac{\delta(x)-\delta(y)}{d(x,y)}.$
It is well-known and easy to prove that unit ball of $\mathcal{F}(M)$ is the closed convex hull of the set of all elementary molecules. Moreover if $Q:\ell_{1}(\widetilde{M})\to\mathcal{F}(M)$ denotes the linear map such that
$Q(e_{(x,y)})=m_{xy},$
then $Q$ extends to an onto norm-one mapping which satisfies
$(\star)\qquad\forall\mu\in\mathcal{F}({M}),\quad\left\|{\mu}\right\|=\inf\left\{{\left\|{a}\right\|_{1}:Qa=\mu}\right\}.$
[…]
We briefly recall the so-called multiplication operators and their preadjoints, the weighted Lipschitz operators. Following [8, Lemma 2.3], let $h:M\to\mathbb{R}$ be a Lipschitz function with bounded support, and let $K\subset M$ contain both the base point and the support of $h$. For $f\in\operatorname{Lip}_{0}(K)$, define
(1) $\mathsf{M}_{h}(f)(x)=\begin{cases}f(x)h(x)&\text{if }x\in K,\\ 0&\text{if }x\notin K.\end{cases}$
Then $\mathsf{M}_{h}$ is a weak-to-weak continuous operator from $\operatorname{Lip}_{0}(K)$ into $\operatorname{Lip}_{0}(M)$, called a multiplication operator. Its preadjoint $T_{h}:\mathcal{F}(M)\to\mathcal{F}(K)$, given by $T_{h}(\delta(x))=h(x)\,\delta(f(x))$, is referred to as a weighted Lipschitz operator. Precise conditions on $h$ and $f$ ensuring boundedness of $\mathsf{M}_{h}$ and $T_{h}$ are characterized in [2].
1.3. Some properties related to weak compactness
A sequence $(x_{n})_{n}$ in a Banach space $X$ is weakly Cauchy if for every $x^{*}\in X^{*}$, the sequence $(\left<{x^{*},x_{n}}\right>)_{n}$ converges. A subset $W$ of a Banach space $X$ is called weakly precompact if every sequence $(x_{n})_{n}\subset W$ admits a weakly Cauchy subsequence. Equivalently, by virtue of Rosenthal’s $\ell_{1}$-theorem, $W$ is weakly precompact if it is bounded and no sequence in $W$ is equivalent to the unit vector basis of $\ell_{1}$. Now a Banach space:
  • has the Schur property if every weakly null sequence is actually norm-null;
  • is weakly sequentially complete (wsc) if every weakly Cauchy sequence is weakly convergent.
Of course, every Schur space is weakly sequentially complete. It is also noteworthy that if $X$ has the Schur property then weak and norm compactness coincide. Also, if $X$ is weakly sequentially complete then weak precompactness and relative weak compactness coincide.
Next, a series $\sum x_{n}$ in a Banach space is weakly unconditionally Cauchy (WUC) if $\sum_{n=1}^{\infty}|x^{*}(x_{n})|<\infty$ for every $x^{*}\in X^{*}$. Now a subset $W\subset X$ is a V*-set if for every WUC series $\sum x_{n}^{*}$ in $X^{*}$ one has
$\lim_{n}\sup_{w\in W}\left|{\left<{x_{n}^{*},w}\right>}\right|=0.$
Observe that every weakly precompact set is a V*-set, but not conversely. In fact, a Banach space
  • has property (V*) if all V*-sets are relatively weakly compact (in that case, they agree with the weakly precompact sets too).
In particular, note that if $X$ has property (V*), then $X$ is weakly sequentially complete. The properties we just mentioned are summarized in the following diagram in terms of implications for closed sets:
compactweakly compactweakly precompactV*-setwscproperty (V*)Schur property
1.4. A brief overview in the framework of Lipschitz free spaces
Let us start with a remark with respect to norm compact sets in Lipschitz free spaces. It is a folklore fact that a set $K$ is relatively compact in a Banach space $X$ if and only if it is bounded and can be approximated by finite dimensional subspaces in the following sense: for every $\varepsilon>0$, there exists a finite dimensional subspace $F\subset X$ such that $K\subset F+\varepsilon B_{X}$. In terms of Lipschitz free spaces, we obtain the next, similar statement:
Proposition 1.1.
Let $M$ be any pointed metric space and let $K\subset\mathcal{F}(M)$ be bounded. Then $K$ is relatively compact if and only if for every $\varepsilon>0$, there exists a finite subset $F\subset M$ such that $K\subset\mathcal{F}(F)+\varepsilon B_{\mathcal{F}(M)}$.
This proposition should be compared to [5, Proposition 3.6] which implies that if $W$ is relatively weakly compact in $\mathcal{F}(M)$, then $W$ has Kalton’s property. That is, for every $\varepsilon>0$ and every $\delta>0$, there exists a finite subset $F\subset M$ such that
$W\subset\mathcal{F}([F]_{\delta})+\varepsilon B_{\mathcal{F}(M)}.$
We emphasize that it is proved in [5] that $W\subset\mathcal{F}(M)$ has Kalton’s property if and only if $W$ is tight, meaning that
$\forall\varepsilon>0,\;\exists K\subset M\text{ compact,}\quad W\subset\mathcal{F}(K)+\varepsilon B_{\mathcal{F}(M)}.$
It is also shown in the same paper that, for a complete metric space $M$, every weakly precompact set $W\subset\mathcal{F}(M)$ is tight. This last result has recently been slightly improved in [9] since the authors prove that V*-sets are also tight. As we already mentioned above, tightness is quite far from characterizing relative weak precompactness in free spaces. Indeed, it is enough to consider any infinite and compact metric space $M$ while setting $W=B_{\mathcal{F}(M)}$.
We conclude the section with some examples of complete metric spaces having at least one property which we introduced in the previous subsection. It is known [4] that $\mathcal{F}(M)$ has the Schur property if and only if $M$ is purely 1-unrectifiable, which means that $M$ does not contain any bi-Lipschitz image of a subset $S$ of $\mathbb{R}$ with positive Lebesgue measure. Classical examples of such metric spaces are the middle-third Cantor set, the Koch curve, or even any $p$-snowflake metric space $(M,d^{p})$ where $p\in(0,1)$. Next, it is proved in [14] that $\mathcal{F}(M)$ is weakly sequentially complete for any $M\subset\mathbb{R}^{n}$. This result is improved in [22] since the authors show that $\mathcal{F}(M)$ has property (V*) for any compact subset $M$ of a superreflexive Banach space. Thanks to the compact reduction principle in [5], $\mathcal{F}(X)$ is weakly sequentially complete for any superreflexive Banach space. Finally, the recent preprint [9] contains many interesting results concerning property (V*) in Lipschitz free spaces. Notably, it is proved that $\mathcal{F}(M)$ has property (V*) when the complete metric space $M$ is locally compact and purely 1-unrectifiable, an $\ell_{p}$ space for $1<p<\infty$, or a Carnot group.
2. Basic properties of uniformly regular sets
For convenience of the reader, we recall the following definitions from the introduction.
Definition 2.1.
A bounded set $W\subset\mathcal{F}(M)$ is:
  • uniformly regular if for every $\varepsilon>0$ and every open set $U\subset M$, there exists a compact subset $K\subset U$ such that $W\subset\mathcal{F}(K\cup U^{c})+\varepsilon B_{\mathcal{F}(M)}$.
  • tight if for every $\varepsilon>0$, there exists a compact subset $K\subset M$ such that $W\subset\mathcal{F}(K)+\varepsilon B_{\mathcal{F}(M)}$.
It is clear from the definition that uniform regularity implies tightness (simply take $U=M$ above), but the converse does not hold in general. Indeed, notice that if $M$ is compact, then any bounded subset of $\mathcal{F}(M)$ is tight but not necessarily uniformly regular:
Remark 2.2.
If $M$ has at least one accumulation point $z$, then $W:=B_{\mathcal{F}(M)}$ is not uniformly regular. To see this, set $U:=M\setminus\left\{{z}\right\}$ and $\varepsilon=\frac{1}{2}$. Let $K\subset U$ be compact. Since $z$ is an accumulation point, we may find $x\in M\setminus(K\cup\left\{{z}\right\})$ such that $\frac{d(x,z)}{d(x,K)}\leqslant 1$. Consider a 1-Lipschitz map $f$ such that $f\mathord{\upharpoonright}_{K\cup\{z\}}=0$ and $\left<{f,m_{xz}}\right>=1$. Then, for every $\gamma\in\mathcal{F}(K\cup\{z\})$
$\|m_{xz}-\gamma\|_{\mathcal{F}(M)}\geqslant\langle f,m_{xz}-\gamma\rangle=\langle f,m_{xz}\rangle=1>\frac{1}{2}.$
This proves that $m_{xz}\not\in\mathcal{F}(K\cup\{z\})+\frac{1}{2}B_{\mathcal{F}(M)}$. Hence $W$ is not uniformly regular as $m_{xz}\in W$.
In contrast, if $M$ has no accumulation point, then compact subsets of $M$ are actually finite and therefore tightness, uniform regularity and (weak) compactness are all equivalent in this case by virtue of Proposition 1.1.
We now continue with a few basic observations concerning uniformly regular sets, which will be useful in the sequel. These statements can be checked directly from the definition.
Proposition 2.3.
  1. (1)
    If $W\subset\mathcal{F}(M)$ is uniformly regular and $S\subset W$ then $S$ is uniformly regular.
  2. (2)
    If $W_{1}$ and $W_{2}$ are uniformly regular in $\mathcal{F}(M)$, then $W_{1}+W_{2}$ is uniformly regular.
  3. (3)
    If $W\subset\mathcal{F}(M)$ is uniformly regular then the convex hull $\operatorname{conv}(W)$ is uniformly regular.
  4. (4)
    “Grothendieck’s criterion”: A set $W\subset\mathcal{F}(M)$ is uniformly regular if and only if for every $\varepsilon>0$, there exists a uniformly regular set $W_{\varepsilon}\subset\mathcal{F}(M)$ such that $W\subset W_{\varepsilon}+\varepsilon B_{\mathcal{F}(M)}$.
Remark 2.4.
Uniformly regular sets are not stable under isomorphisms (even isometric) between Lipschitz-free spaces. Indeed, it is well known that if $C$ denotes the middle-third Cantor set and $M=\mathbb{N}$, then
$\mathcal{F}(C)\equiv\ell_{1}\equiv\mathcal{F}(M).$
However, the forthcoming Proposition 3.3 shows that there exists a uniformly regular set $W\subset\mathcal{F}(C)$ which is not weakly compact, whereas this is impossible in $\mathcal{F}(M)$, as noted in Remark 2.2.
On the other hand, as we will show in the three lemmata below, uniformly regular sets are stable under the following special maps: linearizations of Lipschitz maps, weighted operators, and change of base-point maps.
Lemma 2.5.
Let $f\in\operatorname{Lip}_{0}(M,N)$ and assume that $W\subset\mathcal{F}(M)$ is uniformly regular. Then $\widehat{f}(W)$ is uniformly regular in $\mathcal{F}(N)$.
Proof.
Let $V$ be an open set in $N$ and $\varepsilon>0$. Then $U:=f^{-1}(V)$ is an open set in $M$. Therefore, by assumption, there exists a compact set $K\subset U$ such that $W\subset\mathcal{F}(K\cup(M\setminus U))+\varepsilon B_{\mathcal{F}(M)}$. Since $f$ is continuous, $f(K)$ is a compact subset of $V$.
$\widehat{f}(W)$ $\subset\widehat{f}\left(\mathcal{F}(K\cup(M\setminus U))+\varepsilon B_{\mathcal{F}(M)}\right)$
$\subset\widehat{f}\left(\overline{\operatorname{span}}(K\cup(M\setminus U))\right)+\varepsilon\operatorname{Lip}(f)B_{\mathcal{F}(N)}$
$\subset\overline{\operatorname{span}}(f(K\cup(M\setminus U)))+\varepsilon\operatorname{Lip}(f)B_{\mathcal{F}(N)}$
$=\mathcal{F}(f(K\cup(M\setminus U)))+\varepsilon\operatorname{Lip}(f)B_{\mathcal{F}(N)}$
$\subset\mathcal{F}(f(K)\cup(N\setminus V))+\varepsilon\operatorname{Lip}(f)B_{\mathcal{F}(N)}.$
Lemma 2.6.
Let $N\subset M$ be a closed subset. Let $T:\mathcal{F}(M)\to\mathcal{F}(N)$ be a bounded linear operator such that for every $\mu\in\mathcal{F}(M)$ we have $\operatorname{supp}(T\mu)\subset\operatorname{supp}(\mu)\cap N$. If $W\subset\mathcal{F}(M)$ is uniformly regular then $T(W)$ is uniformly regular.
This lemma applies in particular to weighted operators (see the end of Subsection 1.2).
Proof.
Let $V$ be an open subset of $N$ and let $\varepsilon>0$. There is an open $U\subset M$ such that $V=N\cap U$. Notice that $N\setminus V=N\cap M\setminus U$. By hypothesis, there exists a compact $K\subset U$ such that $W\subset\mathcal{F}(K\cup(M\setminus U))+\varepsilon B_{\mathcal{F}(M)}$. Thus
$T(W)$ $\subset\mathcal{F}(\big(K\cup(M\setminus U)\big)\cap N)+\varepsilon\|T\|B_{\mathcal{F}(N)}$
$\subset\mathcal{F}((K\cap N)\cup(N\setminus V))+\varepsilon\left\|{T}\right\|B_{\mathcal{F}(N)},$
which yields the conclusion. ∎
Lemma 2.7.
Let $0\neq a\in M$. And let $\mathcal{F}_{0}(M)$, resp. $\mathcal{F}_{a}(M)$, be the canonical preduals of $\operatorname{Lip}_{0}(M)$, resp. $\operatorname{Lip}_{a}(M)$. Let $T:\mathcal{F}_{0}(M)\to\mathcal{F}_{a}(M)$ be the canonical isometry such that $T(\delta(x))=\delta(x)-\delta(0)$. Then $W\subset\mathcal{F}_{0}(M)$ is uniformly regular iff $T(W)\subset\mathcal{F}_{a}(M)$ is uniformly regular.
Proof.
Let $U\subset M$ be open set, let $\varepsilon>0$. It is enough to show that there exists $K\subset U$ compact such that $T(W)\subset\mathcal{F}_{a}(K\cup M\setminus U)+\varepsilon B_{\mathcal{F}_{a}(M)}$. First, notice that since $U$ is uniformly regular, there exists a compact $L\subset U$ such that for every $w\in W$ there are $x_{i},y_{i}\in L\cup\left\{{0}\right\}\cup M\setminus U$ and $a_{i}\in\mathbb{R}$, for $i\in\left\{{1,\ldots,n}\right\}$, such that $\left\|{w-\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}}\right\|\leqslant\varepsilon$. We now have
$\left\|{Tw-\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}}\right\|=\left\|{T\left(w-\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\right)}\right\|=\left\|{w-\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}}\right\|\leqslant\varepsilon.$
We can thus conclude by setting $K=L\cup\left\{{0}\right\}$ if $0\in U$ or $K=L$ if $0\in M\setminus U$. ∎
In measure spaces, uniform regularity can be restated equivalently in terms of sequences of pairwise disjoint open sets (see e.g. [3, Theorem 5.3.2]). As the main result in this section, we will now prove two similar equivalences for Lipschitz free spaces.
Theorem 2.8.
Let $M$ be a complete metric space. For a bounded $W\subset\mathcal{F}(M)$, the following are equivalent:
  1. $(i)$
    $W$ is uniformly regular,
  2. $(ii)$
    for every sequence $(U_{n})$ of pairwise disjoint open sets in $M$
    $\lim_{n\to\infty}\sup_{\gamma\in W}\operatorname{dist}(\gamma,\mathcal{F}(U_{n}^{c}))=0,$
  3. $(iii)$
    for every sequence $(f_{n})$ in $B_{\operatorname{Lip}_{0}(M)}$ such that the sets $U_{n}=\left\{{x\in M:f_{n}(x)\neq 0}\right\}$ are pairwise disjoint
    $\lim_{n\to\infty}\sup_{\gamma\in W}\left|{\left<{f_{n},\gamma}\right>}\right|=0.$
Proof.
$(i)\implies(ii)$: We prove the contrapositive. Assume that $(ii)$ fails. Passing to a subsequence if necessary, we may assume that there exist $\varepsilon>0$ and a sequence of pairwise disjoint open sets $(U_{n})$ such that $\sup_{\gamma\in W}\operatorname{dist}(\gamma,\mathcal{F}(U_{n}^{c}))>\varepsilon$. For every $n$, let $\gamma_{n}\in W$ such that $\operatorname{dist}(\gamma_{n},\mathcal{F}(U_{n}^{c}))>\varepsilon$. Set $U:=\bigcup_{n}U_{n}$. Now for any compact $K\subset U$, there exists $N\in\mathbb{N}$ such that $K\subset\bigcup_{n=1}^{N}U_{n}$. In particular $K\cup U^{c}\subset U_{N+1}^{c}$. We deduce that
$\operatorname{dist}(\gamma_{N+1},\mathcal{F}(K\cup U^{c}))\geqslant\operatorname{dist}(\gamma_{N+1},\mathcal{F}(U_{N+1}^{c}))>\varepsilon,$
and hence $W$ is not uniformly regular.
$(ii)\implies(iii)$: Let $\varepsilon>0$ and $(f_{n})\subset B_{\operatorname{Lip}_{0}(M)}$ be such that the sets $U_{n}=\left\{{x\in M:f_{n}(x)\neq 0}\right\}$ are pairwise disjoint and that $\sup_{\gamma\in W}\left|{\left<{\gamma,f_{n}}\right>}\right|>\varepsilon$ for every $n$. For a given $k$, let $\gamma\in W$ be such that $\left|{\left<{\gamma,f_{k}}\right>}\right|>\varepsilon$. Let $\mu\in\mathcal{F}(U_{k}^{c})$. We have $\left\|{\gamma-\mu}\right\|\geqslant\left|{\left<{\gamma-\mu,f_{k}}\right>}\right|=\left|{\left<{\gamma,f_{k}}\right>}\right|>\varepsilon$. It follows that $\sup_{\gamma\in W}\operatorname{dist}(\gamma,\mathcal{F}(U^{c}_{k}))>\varepsilon$ for every $k$.
$(iii)\implies(ii)$: Assume that there exist pairwise disjoint open sets $U_{n}$ and elements $\gamma_{n}\in W$ such that $\operatorname{dist}(\gamma_{n},\mathcal{F}(U_{n}^{c}))>\varepsilon$ for some $\varepsilon>0$. For every $n$, the Hahn-Banach theorem yields $f_{n}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\left<{f_{n},\gamma_{n}}\right>>\varepsilon$ and $f_{n}=0$ on $\mathcal{F}(U_{n}^{c})$, i.e. $f_{n}=0$ on $U_{n}^{c}$. Since $\left<{f_{n},\gamma_{n}}\right>>\varepsilon$, this contradicts $(iii)$.
$(ii)$ and $(iii)\implies(i)$: This implication involves a much longer argument, so we will split the proof into several steps. Assume that $W$ satisfies $(ii)$ and $(iii)$.
Step 1: Given $\varepsilon>0$, there is $r<\infty$ such that $W\subset\mathcal{F}(B(0,r))+\varepsilon B_{\mathcal{F}(M)}$.
We will prove the contrapositive, thus assume that there exists $\varepsilon>0$ such that for every $r>0$ we have $W\not\subset\mathcal{F}(B(0,r))+\varepsilon B_{\mathcal{F}(M)}$. Let $\gamma_{1}\in W$ be such that $\|\gamma_{1}\|>\varepsilon$. By the Hahn-Banach theorem there exists $f_{1}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle f_{1},\gamma_{1}\rangle>\varepsilon$. We may additionally assume that there exists $r_{1}>0$ such that $f_{1}\mathord{\upharpoonright}_{B(0,r_{1})^{c}}=0$, as functions with bounded support are weak dense in $B_{\operatorname{Lip}_{0}(M)}$ (see e.g. [8, Lemma 2.2]). Since $W\not\subset\mathcal{F}(B(0,r_{1}))+\varepsilon B_{\mathcal{F}(M)}$, there exists $\gamma_{2}\in W$ such that $\operatorname{dist}(\gamma_{2},\mathcal{F}(B(0,r_{1}))>\varepsilon$. By the Hahn-Banach theorem there exists $f_{2}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle f_{2},\gamma_{2}\rangle>\varepsilon$ and $f_{2}\mathord{\upharpoonright}_{B(0,r_{1})}=0$. As before, we may additionally assume that there exists $r_{2}>r_{1}$ such that $f_{2}\mathord{\upharpoonright}_{B(0,r_{2})^{c}}=0$. Continuing in such a way we will construct a sequence of functions $(f_{n})\subset B_{\operatorname{Lip}_{0}(M)}$ such that the sets $U_{n}=\left\{{x\in M:f_{n}(x)\neq 0}\right\}$ are pairwise disjoint and $\langle f_{n},\gamma_{n}\rangle>\varepsilon$ for every $n\in\mathbb{N}$, in contradiction with $(iii)$.
Step 2: $W$ is tight. Again, we will prove the contrapositive, thus assume $W$ does not have the Kalton’s property (which is equivalent to tightness [5]). Hence there exist $\varepsilon,\delta>0$ such that for every finite subset $F$ we have $W\not\subset\mathcal{F}([F]_{\delta})+\varepsilon B_{\mathcal{F}(M)}$. By Step 1, there exists $r>\delta/4$ such that
$W\subset\mathcal{F}(B(0,r))+\frac{\varepsilon}{4}B_{\mathcal{F}(M)}.$
Let $\gamma_{1}\in W$ be such that $\|\gamma_{1}\|>\varepsilon$. Let $\mu_{1}\in\mathcal{F}(B(0,r))$ be such that $\|\gamma_{1}-\mu_{1}\|\leq\frac{\varepsilon}{4}$. Then $\|\mu\|>\frac{3\varepsilon}{4}$ and thus there exists $g_{1}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle g_{1},\mu_{1}\rangle>\frac{3\varepsilon}{4}$. We may additionally assume that $g_{1}\mathord{\upharpoonright}_{B(0,2r)^{c}}=0$ by replacing $g_{1}$ with a function $g^{\prime}_{1}$ defined by $g^{\prime}_{1}=g_{1}$ on $B(0,r)$, $g^{\prime}_{1}=0$ on $B(0,2r)^{c}$, and extended to $M$ by McShane-Whitney theorem (see e.g. [25, Theorem 1.33]). Furthermore, $\langle g_{1},\gamma_{1}\rangle>\frac{\varepsilon}{2}$. Pick a finitely supported element $\gamma_{1}^{\prime}\in\mathcal{F}(M)$ such that $\langle g_{1},\gamma_{1}^{\prime}\rangle>\frac{\varepsilon}{2}$ and $\|\gamma_{1}-\gamma_{1}^{\prime}\|<\frac{\varepsilon\delta}{16r}$. Let $F_{1}=\operatorname{supp}(\gamma_{1}^{\prime})\cup\left\{{0}\right\}$ and define $f_{1}\colon[F_{1}]_{\delta/2}^{c}\cup F_{1}\rightarrow\mathbb{R}$ in such way that $f_{1}\mathord{\upharpoonright}_{[F_{1}]_{\delta/2}^{c}}=0$ and $f_{1}\mathord{\upharpoonright}_{F_{1}}=\frac{\delta}{4r}g_{1}\mathord{\upharpoonright}_{F_{1}}$. Since $g_{1}\mathord{\upharpoonright}_{B(0,2r)^{c}}=0$, for every $u\in F_{1}$ we have $\left|{g_{1}(u)}\right|\leq 2r$ and thus $\left|{f_{1}(u)}\right|\leq\frac{\delta}{2}$. Thus, since $\frac{\delta}{4r}\leqslant 1$, the Lipschitz constant of $f_{1}$ is at most 1. Extend $f_{1}$ to $M$ using McShane–Whitney Theorem. Then $f_{1}\in B_{\operatorname{Lip}_{0}(M)}$ and $\langle f_{1},\gamma_{1}^{\prime}\rangle=\frac{\delta}{4r}\langle g_{1},\gamma_{1}^{\prime}\rangle>\frac{\varepsilon\delta}{8r}$, and thus $\langle f_{1},\gamma_{1}\rangle>\frac{\varepsilon\delta}{16r}$.
We will continue recursively. Assume we have found $\gamma_{1},\ldots,\gamma_{n}\in W$, finite sets $F_{1},\ldots,F_{n}\subset M$, and functions $f_{1},\ldots,f_{n}\in B_{\operatorname{Lip}_{0}(M)}$. Now, since $W\not\subset\mathcal{F}\left(\left[\bigcup_{i=1}^{n}F_{i}\right]_{\delta}\right)+\varepsilon B_{\mathcal{F}(M)}$, there exists $\gamma_{n+1}\in W$ be such that
$\operatorname{dist}\left(\gamma_{n+1},\mathcal{F}\left(\left[\bigcup_{i=1}^{n}F_{i}\right]_{\delta}\right)\right)>\varepsilon.$
Let $\mu_{n+1}\in\mathcal{F}(B(0,r))$ be such that $\|\gamma_{n+1}-\mu_{n+1}\|\leq\frac{\varepsilon}{4}$. Then
$\operatorname{dist}\left(\mu_{n+1},\mathcal{F}\left(\left[\bigcup_{i=1}^{n}F_{i}\right]_{\delta}\right)\right)>\frac{3\varepsilon}{4}.$
By the Hahn-Banach theorem, there exists $g_{n+1}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle g_{n+1},\mu_{n+1}\rangle>\frac{3\varepsilon}{4}$ and $g_{n+1}\mathord{\upharpoonright}_{\left[\bigcup_{i=1}^{n}F_{i}\right]_{\delta}}=0$. Since $\operatorname{supp}(\mu)\subset B(0,r)$, we may additionally assume that $g_{n+1}\mathord{\upharpoonright}_{B(0,2r)}=0$ as above. Furthermore, $\langle g_{n+1},\gamma_{n+1}\rangle>\frac{\varepsilon}{2}$. Pick a finitely supported element $\gamma_{n+1}^{\prime}\in\mathcal{F}(M)$ such that $\langle g_{n+1},\gamma_{n+1}^{\prime}\rangle>\frac{\varepsilon}{2}$ and $\|\gamma_{n+1}-\gamma_{n+1}^{\prime}\|<\frac{\varepsilon\delta}{16r}$. Let
$F_{n+1}=\operatorname{supp}(\gamma_{n+1}^{\prime})\setminus\left[\bigcup_{i=1}^{n}F_{i}\right]_{\delta}$
and define $f_{n+1}\colon[F_{n+1}]_{\delta/2}^{c}\cup F_{n+1}\cup\left\{{0}\right\}\rightarrow\mathbb{R}$ in such way that $f_{n+1}=0$ on $F_{n+1}]_{\delta/2}^{c}$ and $f_{n+1}=\frac{\delta}{4r}g_{n+1}$ on $F_{n+1}\cup\left\{{0}\right\}$. For every $u\in F_{n+1}$ we have $\left|{g_{n+1}(u)}\right|\leq 2r$ and thus $\left|{f_{n+1}(u)}\right|\leq\frac{\delta}{2}$. Thus the Lipschitz constant of $f_{n+1}$ is at most 1. Extend $f_{n+1}$ to $M$ using McShane–Whitney Theorem. Then $f_{n+1}\in B_{\operatorname{Lip}_{0}(M)}$ and $\langle f_{n+1},\gamma_{n+1}^{\prime}\rangle=\frac{\delta}{4r}\langle g_{n+1},\gamma_{n+1}^{\prime}\rangle>\frac{\varepsilon\delta}{8r}$ and thus $\langle f_{n+1},\gamma_{n+1}\rangle>\frac{\varepsilon\delta}{16r}$ .
Now we have defined our sequences $(f_{n})\subset B_{\operatorname{Lip}_{0}(M)}$ and $(\gamma_{n})\subset W$ such that $\langle f_{n},\gamma_{n}\rangle>\frac{\varepsilon\delta}{16r}$. Let us check that the sets $U_{n}=\left\{{x\in M:f_{n}(x)\neq 0}\right\}$ are pairwise disjoint. Clearly $U_{n}\subset[F_{n}]_{\delta/2}$. Furthermore, since for every $i<n$ we have $F_{i}]_{\delta}\cap F_{n}=\emptyset$, then also $F_{i}]_{\delta/2}\cap[F_{n}]_{\delta/2}=\emptyset$. Hence the sets $U_{n}$ are pairwise disjoint, and this contradicts $(iii)$.
Step 3: $W$ is uniformly regular. Once again, we work by contradiction. So assume there exist $\varepsilon>0$ and an open set $U\subset M$ such that for every compact $K\subset U$, $W\not\subset\mathcal{F}(K\cup U^{c})+\varepsilon B_{\mathcal{F}(M)}$. Since $W$ is tight by Step 2, there exists a compact set $K\subseteq M$ such that $W\subseteq\mathcal{F}(K)+\frac{\varepsilon}{32}B_{\mathcal{F}(M)}$. Let $K_{0}=\left\{{0}\right\}\cap U$. By assumption there exists $\gamma_{1}\in W$ such that $\operatorname{dist}(\gamma_{1},\mathcal{F}(K_{0}\cup U^{c}))>\varepsilon$. Pick $\mu_{1}\in\mathcal{F}(K)$ finitely supported such that $\|\gamma_{1}-\mu_{1}\|<\frac{\varepsilon}{16}$. Then $\operatorname{dist}(\mu_{1},\mathcal{F}(K_{0}\cup U^{c}))>\frac{\varepsilon}{2}$. Let
$C_{1}$ $:=\operatorname{supp}(\mu_{1})\setminus(K_{0}\cup U^{c}),$
$K_{1}$ $:=\bigcup_{u\in C_{1}}B\left(u,\frac{1}{2}d\left(u,K_{0}\cup U^{c}\right)\right)\cap K,$
$U_{1}$ $:=\bigcup_{u\in C_{1}}B^{O}\left(u,\frac{1}{4}d(u,K_{0}\cup U^{c})\right).$
Clearly $K_{1}$ is compact and $K_{1},U_{1}\subseteq U\setminus K_{0}$. We claim that $\operatorname{dist}(\mu_{1},\mathcal{F}(U_{1}^{c}))>\frac{\varepsilon}{8}$. Indeed, since $\operatorname{dist}(\mu_{1},\mathcal{F}(K_{0}\cup U^{c}))>\frac{\varepsilon}{2}$, by the Hahn-Banach theorem there exists $f_{1}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle f_{1},\mu_{1}\rangle>\frac{\varepsilon}{2}$ and $f_{1}\mathord{\upharpoonright}_{K_{0}\cup U^{c}}=0$. Define $g_{1}\colon U_{1}^{c}\cup C_{1}\rightarrow\mathbb{R}$ in such way that $g_{1}\mathord{\upharpoonright}_{U_{1}^{c}}=0$ and $g_{1}\mathord{\upharpoonright}_{C_{1}}=\frac{1}{4}f_{1}\mathord{\upharpoonright}_{C_{1}}$. For every $u\in C_{1}$ and $v\in U_{1}^{c}$ we have
$|g_{1}(u)-g_{1}(v)|=\frac{1}{4}|f_{1}(u)|\leq\frac{1}{4}d(u,K_{0}\cup U^{c})\leq d(u,v).$
Thus the Lipschitz constant of $g_{1}$ is at most 1. Extend $g_{1}$ to $M$ using McShane–Whitney theorem. Then $g_{1}\in B_{\operatorname{Lip}_{0}(M)}$ and $g_{1}\mathord{\upharpoonright}_{U_{1}^{c}}=0$. Note that if $u\in\operatorname{supp}(\mu_{1})\setminus C_{1}$, then $d(u,K_{0}\cup U^{c})=0$ and thus $g_{1}(u)=0=\frac{1}{4}f_{1}(u)$. Therefore $g_{1}\mathord{\upharpoonright}_{\operatorname{supp}(\mu_{1})}=\frac{1}{4}f_{1}\mathord{\upharpoonright}_{\operatorname{supp}(\mu_{1})}$, which gives us $\langle g_{1},\mu_{1}\rangle=\frac{1}{4}\langle f_{1},\mu_{1}\rangle>\frac{\varepsilon}{8}$. Hence $\operatorname{dist}(\mu_{1},\mathcal{F}(U_{1}^{c}))>\frac{\varepsilon}{8}$, which implies $\operatorname{dist}(\gamma_{1},\mathcal{F}(U_{1}^{c}))>\frac{\varepsilon}{16}$.
Now assume we have defined $\gamma_{1},\ldots,\gamma_{n}\in W$, finite sets $C_{1},\ldots,C_{n}\subseteq K$, compact sets $K_{1},\ldots,K_{n}\subseteq U\cap K$, and pairwise distinct open sets $U_{1},\ldots,U_{n}\subseteq U$ in such a way that: For every $i\in\left\{{1,\ldots,n}\right\}$ we have
$\operatorname{dist}(\gamma_{i},\mathcal{F}(U_{i}^{c}))>\frac{\varepsilon}{16},$
$K_{i}=\bigcup_{u\in C_{i}}B\left(u,\frac{1}{2}d\left(u,\bigcup_{j=0}^{i-1}K_{j}\cup U^{c}\right)\right)\cap K,\text{ and }$
$U_{i}=\bigcup_{u\in C_{i}}B^{O}\left(u,\frac{1}{4}d\left(u,\bigcup_{j=0}^{i-1}K_{j}\cup U^{c}\right)\right).$
The set $\bigcup_{i=0}^{n}K_{i}\subseteq U$ is compact, thus by assumption there exists $\gamma_{n+1}\in W$ such that $\operatorname{dist}\left(\gamma_{n+1},\mathcal{F}\left(\bigcup_{i=0}^{n}K_{i}\cup U^{c}\right)\right)>\varepsilon$. Pick $\mu_{n+1}\in\mathcal{F}(K)$ finitely supported such that $\|\gamma_{n+1}-\mu_{n+1}\|<\varepsilon/16$. Then $\operatorname{dist}(\mu_{n+1},\mathcal{F}(\bigcup_{i=0}^{n}K_{i}\cup U^{c}))>\varepsilon/2$. Let
$C_{n+1}$ $:=\operatorname{supp}(\mu_{n+1})\setminus\left(\bigcup_{i=0}^{n}K_{i}\cup U^{c}\right),$
$K_{n+1}$ $:=\bigcup_{u\in C_{n+1}}B\left(u,\frac{1}{2}d\left(u,\bigcup_{i=0}^{n}K_{i}\cup U^{c}\right)\right)\cap K,$
$U_{n+1}$ $:=\bigcup_{u\in C_{n+1}}B^{O}\left(u,\frac{1}{4}d\left(u,\bigcup_{i=0}^{n}K_{i}\cup U^{c}\right)\right).$
Clearly $K_{n+1}$ is compact and both $K_{n+1}$ and $U_{n+1}$ are subsets of $U\setminus\bigcup_{i=0}^{n}K_{i}$. Similarly as before, one can show that $\operatorname{dist}(\gamma_{n+1},\mathcal{F}(U_{n+1}^{c}))>\frac{\varepsilon}{16}$.
So, the only thing left to prove is $U_{i}\cap U_{n+1}=\emptyset$ for every $i\in\left\{{1,\ldots,n}\right\}$. Assume that there exists $v\in U_{i}\cap U_{n+1}$ for some $i\in\left\{{1,\ldots,n}\right\}$. Then there exists $u_{1}\in C_{i}$ and $u_{2}\in C_{n+1}$ such that $d(u_{1},v)<\frac{1}{4}d(u_{1},\bigcup_{j=0}^{i-1}K_{j}\cup U^{c})$ and $d(u_{2},v)<\frac{1}{4}d(u_{2},\bigcup_{j=0}^{n}K_{j}\cup U^{c})$. Clearly $u_{1}\in K_{i}$, and thus $d(u_{2},\bigcup_{j=0}^{n}K_{j}\cup U^{c})\leq d(u_{1},u_{2}).$ Furthermore, $u_{2}\in K$ and $u_{2}\notin K_{i}$, thus $d(u_{1},u_{2})\geq\frac{1}{2}d(u_{1},\bigcup_{j=0}^{i-1}K_{j}\cup U^{c})$ by the definition of $K_{i}$. Therefore
$d(u_{1},u_{2})$ $\leq d(u_{1},v)+d(u_{2},v)$
$<\frac{1}{4}\left(d\left(u_{1},\bigcup_{j=0}^{i-1}K_{j}\cup U^{c}\right)+d\left(u_{2},\bigcup_{j=0}^{n}K_{j}\cup U^{c}\right)\right)\leqslant d(u_{1},u_{2}).$
This is a clear contradiction. Now we have found a sequence of pairwise disjoint open sets $(U_{n})$ such that
$\sup_{\gamma\in W}\operatorname{dist}(\gamma,\mathcal{F}(U_{n}^{c}))\geqslant\operatorname{dist}(\gamma_{n},\mathcal{F}(U_{n}^{c}))>\frac{\varepsilon}{16},$
which contradicts $(ii)$. This ends the proof. ∎
We conclude this section by showing that the notion of uniform regularity does not depend on the ambient space. Indeed, if $0\in N\subset M$ and $W\subset\mathcal{F}(N)$ is uniformly regular relative to $\mathcal{F}(N)$, then it is clearly uniformly regular relative to $\mathcal{F}(M)$ as well. The converse direction, namely whether $W$ remains uniformly regular relative to $\mathcal{F}(N)$ when it is uniformly regular relative to $\mathcal{F}(M)$, is less immediate. To establish this, we will use the following lemma.
Lemma 2.9.
Let $0\in N\subset M$ be closed. Suppose that $(f_{i})_{i\in I}$ are elements of $B_{\operatorname{Lip}_{0}(N)}$ such that the sets $U_{i}=\left\{{x\in N:f_{i}(x)\neq 0}\right\}$ are pairwise disjoint. Then there exist extensions $\tilde{f}_{i}\in 2B_{\operatorname{Lip}_{0}(M)}$ such that $\tilde{f}_{i}\mathord{\upharpoonright}_{N}=f_{i}$ and the sets $\left\{{x\in M:\tilde{f}_{i}(x)\neq 0}\right\}$ are pairwise disjoint.
Proof.
We may assume that all sets $U_{i}$ are non-empty, otherwise $f_{i}=0$ and we may take $\tilde{f}_{i}=0$. Since they are pairwise disjoint, we may also assume that $N\setminus U_{i}$ are non-empty. Define
$\tilde{U}_{i}=\bigcup_{x\in U_{i}}B^{O}\left({x,\frac{d(x,N\setminus U_{i})}{2}}\right),$
where the balls are taken in $M$. Clearly the sets $(\tilde{U}_{i})_{i\in I}$ are open, and moreover they are pairwise disjoint. Indeed, suppose $p\in\tilde{U}_{i}\cap\tilde{U}_{j}$ for $i\neq j\in I$. Then there exist $x\in U_{i}$, $y\in U_{j}$ such that
$p\in B^{O}\left({x,\frac{d(x,N\setminus U_{i})}{2}}\right)\cap B^{O}\left({y,\frac{d(y,N\setminus U_{j})}{2}}\right)$
therefore
$d(x,y)\leqslant d(x,p)+d(p,y)<\frac{d(x,N\setminus U_{i})}{2}+\frac{d(y,N\setminus U_{j})}{2}\leqslant\frac{d(x,y)}{2}+\frac{d(y,x)}{2},$
a contradiction.
We define $\tilde{f}_{i}$ on $U_{i}\cup M\setminus\tilde{U}_{i}$ as
$\tilde{f}_{i}(x):=\begin{cases}f_{i}(x)&\mbox{ if }x\in U_{i}\\ 0&\mbox{ if }x\in M\setminus\tilde{U}_{i}.\end{cases}$
Notice that $N\setminus U_{i}\subset M\setminus\tilde{U}_{i}$ and $\tilde{f}_{i}\mathord{\upharpoonright}_{N}=f_{i}$. We now check that $\|\tilde{f}_{i}\|_{L}\leqslant 2$. Let $x,y\in U_{i}\cup M\setminus\tilde{U}_{i}$. It is easy to see that we only need to check the situation when $x\in U_{i}$ and $y\in M\setminus(N\cup\tilde{U}_{i})$. Fix $\varepsilon>0$ and $z\in N\setminus U_{i}$ such that $d(x,z)\leqslant(1+\varepsilon)d(x,N\setminus U_{i})$. Then we have:
$\left|{\tilde{f}_{i}(x)-\tilde{f}_{i}(y)}\right|=\left|{f_{i}(x)-f_{i}(z)}\right|\leqslant d(x,z)\leqslant(1+\varepsilon)d(x,N\setminus U_{i})\leqslant 2(1+\varepsilon)d(x,y).$
Letting $\varepsilon\to 0$ yields our claim. Now extend $\tilde{f}_{i}$ to $M$ using the McShane–Whitney theorem. This ends the proof, as $\tilde{f}_{i}$ vanishes outside of $\tilde{U}_{i}$. ∎
Proposition 2.10.
Let $M$ be complete and $N\subset M$ be closed. Then a set $W\subset\mathcal{F}(N)$ is uniformly regular with respect to $\mathcal{F}(M)$ if and only if it is uniformly regular with respect to $\mathcal{F}(N)$.
Proof.
The backward implication is a particular case of Lemma 2.5. For the converse, suppose that $W$ is uniformly regular with $\mathcal{F}(M)$ as the ambient space. We may assume $0\in N$. Let $(f_{n})$ be a sequence in $B_{\operatorname{Lip}_{0}(N)}$ such that the sets $\left\{{x\in M:f_{n}(x)\neq 0}\right\}$ are pairwise disjoint. Use Lemma 2.9 to find a sequence of extensions $(\tilde{f}_{n})$ in $2B_{\operatorname{Lip}_{0}(M)}$ with the same disjointness property. By hypothesis and Theorem 2.8 (iii), $\langle\tilde{f}_{n},\gamma\rangle$ converges to 0 uniformly on $\gamma\in W$. Since $W\subset\mathcal{F}({N})$, we have that $\langle\tilde{f}_{n},\gamma\rangle=\langle\tilde{f}_{n}\mathord{\upharpoonright}_{N},\gamma\rangle=\langle f_{n},\gamma\rangle$ also converges to 0 uniformly. Thus, another application of Theorem 2.8 (iii) shows that $W$ is uniformly regular in $\mathcal{F}(N)$. ∎
3. Uniform regularity versus weak compactness
To start, we derive Theorem A as an immediate corollary of Theorem 2.8.
Corollary 3.1.
Let $M$ be a complete metric space. If $W\subset\mathcal{F}({M})$ is a V*-set, in particular if it is weakly compact, then it is uniformly regular.
Proof.
We show that $W$ satisfies condition $(iii)$ in Theorem 2.8. Let $(f_{n})$ be a sequence in $B_{\operatorname{Lip}_{0}(M)}$ such that the sets $U_{n}=\left\{{x\in M:f_{n}(x)\neq 0}\right\}$ are pairwise disjoint. Then $\sum_{n}f_{n}$ is a WUC series in $\operatorname{Lip}_{0}(M)$ by [6, Lemma 6] (see also the proof of [13, Lemma 1.5]). Since $W$ is a V*-set, we have $\lim_{n}\sup_{\gamma\in W}\left|{\left<{f_{n},\gamma}\right>}\right|=0$ as required. ∎
This argument highlights nicely the difference between uniformly regular sets and V*-sets: $W$ is a V*-set if $\lim_{n}\sup_{\gamma\in W}\left|{\left<{f_{k},\gamma}\right>}\right|=0$ for all WUC series $\sum_{n}f_{n}$ in $\operatorname{Lip}_{0}(M)$, while $W$ is uniformly regular if the same holds for WUC series which moreover satisfy that the sets $\left\{{f_{n}\neq 0}\right\}$ are pairwise disjoint.
This points to the fact that V*-sets and uniformly regular sets may not coincide in general. Although it is not difficult to find a set $W\subset\mathcal{F}(M)$ which is tight but not V* (or not weakly compact), it is more challenging to identify a counterexample in the case of uniform regularity. Our next aim is to address this challenge. To begin, we present a technical lemma.
Lemma 3.2.
Let $M$ be a metric space, and let $U\subseteq M$ be open. If $\varepsilon\in(0,1)$ and $\gamma=\frac{1}{n}\sum_{i=1}^{n}m_{u_{i}v_{i}}\in\mathcal{F}({M})$ are such that $\operatorname{dist}(\gamma,\mathcal{F}({U^{c}}))>\varepsilon$, then
$\big|\big\{i\in\{1,\ldots,n\}:B(u_{i},r_{i})\subseteq U\text{ or }B(v_{i},r_{i})\subseteq U\big\}\big|>\frac{n\varepsilon}{2},$
where $r_{i}=\frac{\varepsilon}{8}d(u_{i},v_{i})$, $i\in\{1,\ldots,n\}$.
Proof.
Denote
$I=\big\{i\in\{1,\ldots,n\}:B(u_{i},r_{i})\subseteq U\text{ or }B(v_{i},r_{i})\subseteq U\big\}.$
Note that for every $i\in\{1,\ldots,n\}\setminus I$, there exist $x_{i}\in B(u_{i},r_{i})\cap U^{c}$ and $y_{i}\in B(v_{i},r_{i})\cap U^{c}$. Thus, by [23, Lemma 1.2]:
$\|m_{u_{i}v_{i}}-m_{x_{i}y_{i}}\|$ $=\frac{d(u_{i},x_{i})+d(v_{i},y_{i})+|d(u_{i},v_{i})-d(x_{i},y_{i})|}{\max\{d(u_{i},v_{i}),d(x_{i},y_{i})\}}$
$\leq\frac{d(u_{i},x_{i})+d(v_{i},y_{i})+d(u_{i},x_{i})+d(v_{i},y_{i})}{d(u_{i},v_{i})}$
$\leq\frac{4r_{i}}{d(u_{i},v_{i})}=\frac{\varepsilon}{2}.$
Let $\nu=\frac{1}{n}\sum_{i\notin I}m_{x_{i}y_{i}}$. Clearly $\nu\in\mathcal{F}({U^{c}})$ and
$\varepsilon<\|\gamma-\nu\|\leq\frac{1}{n}\sum_{i\in I}\|m_{u_{i}v_{i}}\|+\frac{1}{n}\sum_{i\notin I}\|m_{u_{i}v_{i}}-m_{x_{i}y_{i}}\|\leq\frac{|I|}{n}+\frac{\varepsilon}{2}$
therefore $|I|>n\varepsilon/2$. ∎
Proposition 3.3.
Let $M$ be a complete metric space containing a non-empty perfect subset. Then $\mathcal{F}(M)$ contains a subset that is uniformly regular but not a V*-set.
Proof.
Without loss of generality, we may assume that $M$ is a complete perfect metric space. There exists a non-empty closed perfect subset $C\subset M$ which can be written as $C=\bigcap_{n}\bigcup S_{n}$ where $S_{n}=\left\{{A_{s}:s\in\left\{{0,1}\right\}^{n}}\right\}$ and moreover we have for all $s\in\left\{{0,1}\right\}^{<\mathbb{N}}$ and all $i\in\left\{{0,1}\right\}$,
  1. (1)
    $A_{s}$ is perfect, closed, and non-empty,
  2. (2)
    $A_{s^{\smallfrown}i}\subset A_{s}$,
  3. (3)
    $\operatorname{diam}(A_{s^{\smallfrown}i})\leqslant\frac{1}{3}\operatorname{diam}(A_{s})$,
  4. (4)
    $d(A_{s^{\smallfrown}0},A_{s^{\smallfrown}1})\geqslant\frac{1}{3}\operatorname{diam}(A_{s})$.
Such a system indeed exists. To see this, recall that the closure of every open subset of a perfect set is perfect. Set $A_{\emptyset}:=\overline{B^{O}(0,\frac{1}{2})}$ and when $A_{s}$ has been defined take $x_{0},x_{1}\in A_{s}$ such that $d(x_{0},x_{1})>\frac{7}{9}\operatorname{diam}(A_{s})$, then define $A_{s^{\smallfrown}i}:=\overline{B^{O}(x_{i},\frac{1}{9}\operatorname{diam}(A_{s}))\cap A_{s}}$. The required properties follow easily.
Given $C$ and $\left\{{S_{n}}\right\}_{n=0}^{\infty}$ as above, we assume without loss of generality that $\operatorname{diam}(A_{\emptyset})=1$. Now we define a system $A^{\prime}_{s}$, $s\in\left\{{0,1}\right\}^{<\mathbb{N}}$ of non-degenerate closed subintervals of $[0,1]$ satisfying properties (1)-(4) above and such that $\operatorname{diam}(A^{\prime}_{s})=\operatorname{diam}(A_{s})$. First, let $A_{\emptyset}^{\prime}=[0,1]$. Next, assume that $A_{s}^{\prime}$ has been defined and choose $A_{s^{\smallfrown}0}^{\prime}$ and $A_{s^{\smallfrown}1}^{\prime}$ to be placed on the left and right sides of $A^{\prime}_{s}$, respectively, i.e.
$A^{\prime}_{s^{\smallfrown}0}$ $:=[\min A_{s}^{\prime},\min A_{s}^{\prime}+\operatorname{diam}(A_{s^{\smallfrown}0})]$
$A^{\prime}_{s^{\smallfrown}1}$ $:=[\max A^{\prime}_{s}-\operatorname{diam}(A_{s^{\smallfrown}1}),\max A^{\prime}_{s}].$
It is clear that they satisfy the required properties as well. For further reference, let us denote $u_{s^{\smallfrown}0}^{\prime}:=\max A_{s^{\smallfrown}0}^{\prime}$ and $u_{s^{\smallfrown}1}^{\prime}:=\min A_{s^{\smallfrown}1}^{\prime}$, and notice that the interval $(u^{\prime}_{s^{\smallfrown}0},u^{\prime}_{s^{\smallfrown}1})$ is contained in $A_{s}^{\prime}$ but intersects neither $A^{\prime}_{s^{\smallfrown}0}$ nor $A^{\prime}_{s^{\smallfrown}1}$.
We define the generalized Cantor set $C^{\prime}:=\bigcap_{n}\bigcup S_{n}^{\prime}$ where $S^{\prime}_{n}=\left\{{A_{s}^{\prime}:s\in\left\{{0,1}\right\}^{n}}\right\}$. To simplify notation, for $s=(s_{k})\in\left\{{0,1}\right\}^{\mathbb{N}}$ we write $s_{\mathord{\upharpoonright}n}:=(s_{1},\ldots,s_{n})$. It is well known, and easily verified, that the above properties imply that for every $s=(s_{n})\in\{0,1\}^{\mathbb{N}}$ there exist points $x_{s}$ and $x_{s}^{\prime}$ such that
$\{x_{s}\}=\bigcap_{n}A_{s_{\mathord{\upharpoonright}n}},\qquad\{x_{s}^{\prime}\}=\bigcap_{n}A_{s_{\mathord{\upharpoonright}n}}^{\prime},$
and moreover this completely describes all the points of $C$ and $C^{\prime}$, respectively. We claim that the map $\varphi:x_{s}\mapsto x_{s}^{\prime}$ is bi-Lipschitz. Indeed, let $x_{s}\neq x_{t}\in C$ be given and let $n\in\mathbb{N}$ be the smallest integer such that $s_{n+1}\neq t_{n+1}$. Then $x_{s}^{\prime},x_{t}^{\prime}\in A_{s_{\mathord{\upharpoonright}n}}^{\prime}$ and therefore $\left|{x_{s}^{\prime}-x_{t}^{\prime}}\right|\leqslant\operatorname{diam}(A^{\prime}_{s_{\mathord{\upharpoonright}n}})=\operatorname{diam}(A_{s_{\mathord{\upharpoonright}n}})$. On the other hand, $x_{s}\in A_{s_{\mathord{\upharpoonright}n}^{\smallfrown}i}$ and $x_{t}\in A_{s_{\mathord{\upharpoonright}n}^{\smallfrown}j}$ with $i\neq j$, so that $d(x_{s},x_{t})\geqslant\frac{1}{3}\,\mathrm{diam}(A_{s_{\mathord{\upharpoonright}n}})$. Therefore,
$|x_{s}^{\prime}-x_{t}^{\prime}|\leqslant 3\,d(x_{s},x_{t}).$
A Lipschitz estimate for the inverse is obtained by interchanging the roles of $C$ and $C^{\prime}$. Notice also that the points $u_{s}^{\prime}$ defined above belong to the generalized Cantor set $C^{\prime}$ and that the corresponding points $u_{s}=\varphi^{-1}(u_{s}^{\prime})$ satisfy $u_{s}\in C\cap A_{s}$.
Let now
$\gamma_{n}:=\frac{1}{2^{n-1}}\sum_{s\in\left\{{0,1}\right\}^{n-1}}m_{u_{s^{\smallfrown}1}u_{s^{\smallfrown}0}}\in\mathcal{F}(C)\,\mbox{ and }\,\gamma_{n}^{\prime}=\widehat{\varphi}(\gamma_{n})\in\mathcal{F}(C^{\prime}).$
We denote $W:=\left\{{\gamma_{n}:n\in\mathbb{N}}\right\}$ and $W^{\prime}=\widehat{\varphi}(W)$. Assume towards a contradiction that $W$ is a V*-set. Consider a Lipschitz extension $\psi:M\to[0,1]$ of $\varphi$. Then $W^{\prime}=\widehat{\varphi}(W)=\widehat{\psi}(W)$ (see for example [8, Proposition 2.6]). Since V*-sets are stable under bounded linear operators, $W^{\prime}$ is a V*-set in $\mathcal{F}([0,1])=L_{1}$. Since $L_{1}$ enjoys property (V*), $W$ must be relatively weakly compact. But $(\gamma_{n}^{\prime})_{n}$ is equivalent to the $\ell_{1}$-basis as we will show now. To do this, we use the usual isometry $T:\delta(x)\in\mathcal{F}([0,1])\mapsto{\mathbf{1}}_{{[0,x]}}\in L^{1}([0,1])$. Under this transformation, we have
$f_{n}:=T(\gamma_{n}^{\prime})=\frac{1}{2^{n-1}}\sum_{s\in\left\{{0,1}\right\}^{n-1}}\frac{{\mathbf{1}}_{{(u_{s^{\smallfrown}0}^{\prime}u_{s^{\smallfrown}1}^{\prime})}}}{d(u_{s^{\smallfrown}1},u_{s^{\smallfrown}0})}.$
By the choice of points $(u^{\prime}_{s})$, it is evident that the maps $f_{n}$ have disjoint support. Moreover, since $\varphi$ is bi-Lipschitz and since the sets $A_{s}^{\prime}$ for $s\in\left\{{0,1}\right\}^{n-1}$ are pairwise disjoint, we have that $\frac{1}{3}\leqslant\left\|{f_{n}}\right\|\leqslant 3$, indicating that $(f_{n})_{n}$ is equivalent to the $\ell_{1}$-basis (see e.g. [3, Lemma 5.1.1]). It follows that $(\gamma_{n})$ is also equivalent to the $\ell_{1}$-basis, as $\gamma_{n}^{\prime}=\widehat{\varphi}(\gamma_{n})$ and $\widehat{\varphi}$ is an isomorphism. Hence $W$ is not a V*-set.
It remains to be shown that the set $W=\left\{{\gamma_{n}:n\in\mathbb{N}}\right\}$ is uniformly regular. Let $(U_{n})$ be a sequence of pairwise disjoint open subsets of $M$. In order to reach contradiction via Theorem 2.8 we assume that
$\limsup_{k\to\infty}\sup_{n\in\mathbb{N}}\operatorname{dist}(\gamma_{n},\mathcal{F}(U^{c}_{k}))>\varepsilon$
for some $\varepsilon>0$. Let $l\in\mathbb{N}$ be such that $3^{-l}<\frac{\varepsilon}{8\cdot 3}$. Since finite subsets of $\mathcal{F}(M)$ are uniformly regular, we may choose sequences $(n_{i})\subset\mathbb{N}$ and $(k_{i})\subset\mathbb{N}$ such that $n_{i+1}>n_{i}+l$ and $k_{i+1}>k_{i}$ and such that $\operatorname{dist}(\gamma_{n_{i}},\mathcal{F}(U^{c}_{k_{i}}))>\varepsilon$. To simplify the notation we write, as we may, $\operatorname{dist}(\gamma_{n_{k}},\mathcal{F}(U^{c}_{k}))>\varepsilon$ for every $k\in\mathbb{N}$.
We now apply Lemma 3.2 (and the property (4) of the Cantor scheme above) to obtain for every $k\in\mathbb{N}$ that
$\left|{\left\{{s\in\left\{{0,1}\right\}^{n_{k}-1}:B(u_{s^{\smallfrown}0},r_{s})\subset U_{k}\mbox{ or }B(u_{s^{\smallfrown}1},r_{s})\subset U_{k}}\right\}}\right|>2^{n_{k}-2}\cdot\varepsilon$
where $r_{s}=\frac{\varepsilon}{8}\frac{\operatorname{diam}(A_{s})}{3}$. Notice that, since $3^{-l}<\frac{\varepsilon}{8\cdot 3}$, we have for every $s\in\left\{{0,1}\right\}^{n_{k}-1}$ and every $t\in\left\{{0,1}\right\}^{l}$ that
$\operatorname{diam}(A_{s^{\smallfrown}t})\leqslant\left(\frac{1}{3}\right)^{l}\operatorname{diam}(A_{s})\leqslant\frac{\varepsilon}{8}\frac{\operatorname{diam}(A_{s})}{3}=r_{s}.$
It follows that $A_{s^{\smallfrown}t}\subset U_{k}$ when $B(u_{s^{\smallfrown}i},r_{s})\subset U_{k}$ and $u_{s^{\smallfrown}i}\in A_{s^{\smallfrown}t}$. That is, for every $k$, at least $2^{n_{k}-2}\cdot\varepsilon$ different sets $A_{s}\in S_{n_{k}-1+l}$ are included in $U_{k}$. Further, because of the splitting-in-two-property of the Cantor scheme, for every $m\in\mathbb{N}$, and every $k$, we will have at least $2^{m}\cdot 2^{n_{k}-2}\cdot\varepsilon$ different sets $A_{s}\in S_{n_{k}-1+l+m}$ that are included in $U_{k}$. Let $K\in\mathbb{N}$. For every $k=1,\ldots,K$ we choose $m$ above as $n_{K}-n_{k}$. Thus at least $2^{n_{K}-n_{k}}\cdot 2^{n_{k}-2}\cdot\varepsilon=2^{n_{K}-2}\cdot\varepsilon$ different sets $A_{s}\in S_{n_{k}-1+l+n_{K}-n_{k}}=S_{n_{K}-1+l}$ are included in $U_{k}$. Thus, since the sets $(U_{k})$ are pairwise disjoint, $K\cdot 2^{n_{K}-2}\cdot\varepsilon$ different members of $S_{n_{K}-1+l}$ are included in $\bigcup_{k=1}^{K}U_{k}$. But the cardinality of $S_{n_{K}-1+l}$ is $2^{n_{K}-1+l}$. Therefore, if $K$ is chosen large enough, the fact that $(U_{k})$ are pairwise disjoint leads to a contradiction. ∎
As the byproduct of the above proof, we can state the following lemma of independent interest, which is probably well known but we could not locate in the literature.
Lemma 3.4.
Let $M$ be a perfect metric space. Then $M$ contains a bi-Lipschitz copy of a perfect subset of $\mathbb{R}$.
Definition 3.5.
We will say that $M$ has property:
[…]
Next, for a metric space $M$, we define its subset
$M_{k}=\left\{{x\in M:d(x,0)\in[2^{k},2^{k+2}]}\right\}\cup\{0\}$
for every $k\in\mathbb{Z}$. The well-known Kalton decomposition [21, Proposition 4.3] asserts that there exist Lipschitz functions $\varphi_{n}:M\to\mathbb{R}$ with $\operatorname{supp}(\varphi_{n})\subset M_{n}$ such that the pre-adjoints $T_{n}=T_{\varphi_{n}}$ of the associated multiplication operators $\mathsf{M}_{\varphi_{n}}:\operatorname{Lip}_{0}(M_{n})\to\operatorname{Lip}_{0}(M)$ satisfy that
$T:=\bigoplus_{n\in\mathbb{Z}}T_{n}:\mathcal{F}(M)\to X,\qquad\text{where }X=\left(\sum_{n\in\mathbb{Z}}\mathcal{F}(M_{n})\right)_{\ell_{1}},$
is a bounded operator with complemented range. More precisely, we have $S\circ T=\mathrm{Id}_{\mathcal{F}(M)}$, where $S:\left(\sum_{n\in\mathbb{Z}}\mathcal{F}(M_{n})\right)_{\ell_{1}}\to\mathcal{F}(M)$ is the sum of the coordinates. The following propositions will be our main workhorses.
Proposition 3.6.
Let $M$ be a pointed metric space, let $W\subset\mathcal{F}(M)$ be uniformly regular. Then for every $\varepsilon>0$ there exists a finite set $F\subset\mathbb{Z}$ such that
$T(W)\subset T_{F}(W)+\varepsilon B_{X}$
where $T_{F}=\bigoplus_{n\in F}T_{n}$.
Proof.
Let $U=\bigcup_{n\in\mathbb{Z}}M_{n}=M\setminus\{0\}$ and let $\varepsilon>0$ be fixed. There exists a compact $K\subset U$ such that $W\subset\mathcal{F}(K\cup\left\{{0}\right\})+\frac{\varepsilon}{2\left\|{T}\right\|}B_{\mathcal{F}(M)}$. Since $\inf_{x\in K}d(0,x)>0$ and $\sup_{x\in K}d(0,x)<\infty$, there is a finite set $F\subset\mathbb{Z}$ such that $\varphi_{n}=0$ on $K$ whenever $n\notin F$. Let $w\in W$ and $\tilde{w}\in\mathcal{F}(K\cup\left\{{0}\right\})$ such that $\left\|{w-\tilde{w}}\right\|\leqslant\frac{\varepsilon}{2\left\|{T}\right\|}$. Then
$\left\|{Tw-T_{F}w}\right\|=\left\|{Tw-T\tilde{w}+T_{F}\tilde{w}-T_{F}w}\right\|\leqslant 2\left\|{T}\right\|\left\|{w-\tilde{w}}\right\|\leqslant\varepsilon$
as desired. ∎
Proposition 3.7.
[…]
Proof.
[…]
$\sup_{w\in W}\left|{\left<{T_{F}w,f_{k}}\right>}\right|\leqslant\sum_{n\in F}\sup_{w\in W}\left|{\left<{T_{n}w,f_{k,n}}\right>}\right|\to 0$
as $k\to\infty$. Thus, by the Grothendieck criterion for V*-sets (see e.g. [9, Fact 1.6]), $T(W)$ is a V*-set. Finally, since the image of V*-set by a bounded operator is a V*-set, $W=S\circ T(W)$ is a V*-set. ∎
In the next statement, if $\alpha$ is an ordinal then $M^{(\alpha)}$ denotes the Cantor-Bendixson derived set of order $\alpha$.
Proposition 3.8.
[…]
Proof.
The proof proceeds by transfinite induction on $\alpha$. Let us denote:
[…]
If $M$ is finite then $\dim\mathcal{F}(M)<\infty$, and so every bounded set is relatively compact.
[…]
Theorem 3.9.
Let $M$ be a complete metric space. Then the following assertions are equivalent:
  1. $(i)$
    $M$ is scattered.
  2. $(ii)$
[…]
  3. $(iii)$
[…]
  4. $(iv)$
[…]
  5. $(v)$
[…]
Proof.
First, observe that the chain of implications $(ii)\Rightarrow(iii)\Rightarrow(iv)\Rightarrow(v)$ is immediate. On the other hand, if $M$ is not scattered, then it contains a non-empty perfect subset, and we can apply Proposition 3.3 to obtain $(v)\Rightarrow(i)$. Consequently, it remains only to prove $(i)\Rightarrow(ii)$. Note that, under the additional assumption that $M$ is compact, scatteredness is equivalent to countability, and therefore this implication follows from Proposition 3.8. To establish it in the general case, we proceed by means of a compact reduction argument. Let $M$ be scattered. Assume that $W\subset\mathcal{F}(M)$ is uniformly regular and fix $\varepsilon>0$. Then it is tight and [5, Theorem 2.3] implies that there exist a compact $K\subset M$ and a sequence of weighted operators $T_{n}:\mathcal{F}(M)\to\mathcal{F}(M)$ which converges uniformly on $W$ to a mapping $T:W\to\mathcal{F}(K)$, and moreover $W\subset T(W)+\varepsilon B_{\mathcal{F}(M)}$. So, by Lemma 2.6, the sets $T_{n}(W)$ are uniformly regular. By the uniform convergence the set $T(W)$ is arbitrarily close to these sets. Proposition 2.3 (4) (“Grothendieck’s criterion”) thus shows that $T(W)$ is uniformly regular with respect to $\mathcal{F}(M)$. Proposition 2.10 further shows that $T(W)$ is uniformly regular with respect to $\mathcal{F}(K)$. The compact case implies that $T(W)$ is relatively compact in $\mathcal{F}(K)$, hence in $\mathcal{F}(M)$. Recalling that $W\subset T(W)+\varepsilon B_{\mathcal{F}(M)}$, “Grothendieck’s criterion for compacts” implies that $W$ is relatively compact itself. This shows $(i)\implies(ii)$. ∎
Corrollary C is now obtained as a consequence of Corollaries 3.1 and 3.9. This previously unknown result improves [9, Corollary 2.10].
Corollary 3.10.
If $M$ is complete and scattered, then $\mathcal{F}(M)$ has property (V*).
To conclude this section, we introduce another class of subsets of Lipschitz free spaces for which uniform regularity serves as a characterization of relative compactness. If $k\in\mathbb{N}$, let $\mathcal{FS}_{k}(M)$ denote the subset of elements in $\mathcal{F}(M)$ whose support contains at most $k$ elements. For sequences in $\mathcal{FS}_{k}(M)$, norm and weak convergence are actually equivalent [1, Theorem C].
Proposition 3.11.
Let $M$ be a complete metric space and let $W\subset\mathcal{FS}_{k}(M)$ be bounded. Then $W$ is relatively compact if and only if it is uniformly regular.
Proof.
By Corollary 3.1, it suffices to show that if $W\subset\mathcal{FS}_{k}(M)$ is uniformly regular, then $W$ is relatively compact. Let $(\gamma_{n})_{n}$ be a sequence in $W$. By assumption, each element $\gamma_{n}$ can be written as
$\gamma_{n}=\sum_{i=1}^{k}a_{i}^{n}\,\delta(x_{i}^{n}),\qquad\text{for some $a_{i}^{n}\in\mathbb{R}$ and $x_{i}^{n}\in M$.}$
By passing to a subsequence (which we continue to denote $(\gamma_{n})_{n}$ for simplicity), we may assume that for each $i\in\{1,\ldots,k\}$, the sequence $(x_{i}^{n})_{n}$ is either Cauchy (hence convergent) or uniformly discrete. Let
$S:=\overline{\{x_{i}^{n}:n\in\mathbb{N},\ 1\leq i\leq k\}}\subset M.$
Then $S$ is a complete subset of $M$ with at most $k$ accumulation points, hence scattered. Moreover, since every $\gamma_{n}$ is supported on $S$, we have $\{\gamma_{n}:n\in\mathbb{N}\}\subset\mathcal{F}(S)$. By Proposition 2.3 (1) together with Proposition 2.10, the set $\{\gamma_{n}:n\in\mathbb{N}\}$ is uniformly regular in $\mathcal{F}(S)$. Applying Corollary 3.9, we conclude that $(\gamma_{n})_{n}$ admits a convergent subsequence. Hence $W$ is relatively compact, as desired. ∎
4. Some insights from the case of real trees
An $\mathbb{R}$-tree is an arc-connected metric space $(T,d)$ with the property that there is a unique arc connecting any pair of points $x\neq y\in T$ and it is moreover isometric to the real segment $0,d(x,y)]\subset\mathbb{R}$. Such an arc is denoted by $[x,y]$ and is called a segment of $T$. We will also use the obvious notation $(x,y)=[x,y]\setminus\{x,y\}$, and sets of this type will sometimes be called intervals. On every real tree $T$, there is a canonical measure $\lambda_{T}$, called the length measure, whose restriction to every interval is equal to (the pushforward of) Lebesgue measure. The operator $I:\mathcal{F}(T)\to L^{1}(\lambda_{T})$ given by $I\delta(x)={\mathbf{1}}_{{[0,x]}}$ is a surjective isometry. If $0\in T^{\prime}\subset T$ is a subtree, then $\lambda_{T^{\prime}}$ is just the restriction of $\lambda_{T}$ to $T^{\prime}$, and $I$ restricts to an isometry between $\mathcal{F}(T^{\prime})$ and $L^{1}(\lambda_{T^{\prime}})=L^{1}(T^{\prime},\lambda_{T})$. Note also that if a family of segments $x_{i},y_{i}]\subset T$ intersect pairwise trivially (at one point at most) then so do the supports of $I(m_{x_{i}y_{i}})$, and therefore the molecules $m_{x_{i}y_{i}}\in\mathcal{F}(T)$ are an isometric $\ell_{1}$ basis (of appropriate cardinality). For more information on $\mathbb{R}$-trees and their Lipschitz free spaces, we refer the reader to Godard’s article [16].
In this section, we will use Godard’s isometry $G:\mathcal{F}(T)\to L^{1}(\lambda_{T})$ to describe the relatively weakly compact sets in $\mathcal{F}(T)$. To do so, it is necessary to recall first the notion of equi-integrability in $L^{1}$-spaces. Let $(\Omega,\mathcal{A},\mu)$ be a measure space (not necessarily finite or even $\sigma$-finite). Then we say that a set $H\subset L^{1}(\mu)$ is equi-integrable if it is bounded and satisfies the following conditions:
  1. $(a)$
    For any $\varepsilon>0$, there exists $\delta>0$ such that for every $B\in\mathcal{A}$, $\mu(B)<\delta$ implies that $\int_{B}|f|\,d\mu<\varepsilon$ for all $f\in H$.
  2. $(b)$
    For any $\varepsilon>0$, there exists $B\in\mathcal{A}$ with $\mu(B)<\infty$ s.t. $\int_{\Omega\setminus B}|f|\,d\mu<\varepsilon$ for all $f\in H$.
Notice that when $\mu$ is finite, condition (a) implies condition (b). According to the Dunford–Pettis theorem, a set $H$ is relatively weakly compact in $L^{1}(\mu)$ if and only if it is equi-integrable (see e.g. [24, Theorem 15.4]). It is worth noting that, since $L^{1}$-spaces are weakly sequentially complete (see [3, Theorem 5.2.9]), weakly precompact sets coincide with relatively weakly compact sets. In what follows, we will transfer the equi-integrability conditions through the isometry $L^{1}(\lambda_{T})\to\mathcal{F}(T)$ and describe the resulting properties in $\mathcal{F}(T)$.
We shall examine the condition $(a)$ in the proposition below. For convenience, we remind the reader that a convex series of molecules is a series of the form $\gamma=\sum_{i=1}^{\infty}a_{i}m_{x_{i}y_{i}}$ with moreover $\|\gamma\|=\sum_{i=1}^{\infty}|a_{i}|$. If the sum is finite, we will say convex sum of molecules instead. Convex sums and convex series of molecules should not be confused with convex combinations of molecules. We will show that property $(a)$ is related to the distribution of the weights $|a_{i}|$ with respect to the distances $d(x_{i},y_{i})$ between pairs of points defining the molecules.
Proposition 4.1.
Let $T$ be a $\mathbb{R}$-tree. We denote by $G:\mathcal{F}(T)\to L^{1}(\lambda_{T})$ Godard’s isometry. For a bounded $W\subset\mathcal{F}(T)$, the following are equivalent:
  1. $(i)$
    $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every measurable $B\subset T$,
    $\lambda_{T}(B)<\delta\implies\int_{B}|Gw|\,d\lambda_{T}<\varepsilon.$
  2. $(ii)$
    $\forall\varepsilon>0$, $\exists\delta>0$ such that for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$ $\forall I\subset\{1,\ldots,n\}$,
    $\sum_{i\in I}d(x_{i},y_{i})<\delta\implies\sum_{i\in I}\left|{a_{i}}\right|\leqslant\varepsilon+\operatorname{dist}(\gamma,W).$
  3. $(iii)$
    $\exists C>0$, $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$ with $(x_{i})_{i=1}^{n}\cup(y_{i})_{i=1}^{n}\subset\operatorname{supp}(w)\cup\left\{{0}\right\}$, $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\forall I\subset\{1,\ldots,n\}$,
    $\sum_{I}d(x_{i},y_{i})<\delta\implies\sum_{I}|a_{i}|\leq\varepsilon+C\varepsilon.$
In order to prove $(i)\implies(ii)$ we need the following lemma. Observe that for every $x\neq y\in T$ we have $G(m_{xy})=\frac{1}{d(x,y)}({\mathbf{1}}_{{[x,b]}}-{\mathbf{1}}_{{[b,y]}})$ where $b\in T$ is the unique point in $[x,y]$ such that $d(0,b)=d(0,[x,y])$.
Lemma 4.2.
Let $\gamma=\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$ be a convex sum of molecules. Let $I\subset\{1,\ldots,n\}$. Then there exists $f_{I}\in L_{\infty}(\lambda_{T})$, $\|f_{I}\|=1$ such that $\left<{f_{I},G\gamma}\right>\geqslant\sum_{i\in I}|a_{i}|$ and $\operatorname{supp}(f_{I})\subset\bigcup_{i\in I}[x_{i},y_{i}]$.
Proof.
We may assume that $a_{i}\neq 0$ for every $i\in\left\{{1,\ldots,n}\right\}$. Then, notice that $\operatorname{supp}(G(a_{i}m_{x_{i}y_{i}}))=[x_{i},y_{i}]$. Since $\gamma$ is a convex sum of molecules, [11, Theorem 2.4] yields that $\lambda_{T}([x_{i},y_{i}]\cap[x_{j},y_{j}])>0$ implies that $\mathrm{sign}(G(a_{i}m_{x_{i}y_{i}}))(t)=\mathrm{sign}(G(a_{j}m_{x_{j}y_{j}}))(t)$ for almost every $t\in[x_{i},y_{i}]\cap[x_{j},y_{j}]$. We can thus define $f_{I}(t):=\mathrm{sign}(G(a_{i}m_{x_{i}y_{i}}))$ whenever $t\in[x_{i},y_{i}]$ for some $i\in I$, and $f_{I}(t):=0$ otherwise. Now
$\left<{f_{I},G\gamma}\right>=\sum_{i\in I}\left|{a_{i}}\right|+\sum_{i\notin I}\left|{a_{i}}\right|\sum_{j\in I}\lambda_{T}([x_{i},y_{i}]\cap[x_{j},y_{j}]).$
Proof of the downward pointing implications of Proposition 4.1.
(i) $\implies$ (ii): Let us fix $\varepsilon>0$ and let $\delta>0$ be obtained from (i). Let $\gamma\in F(M)$ be a convex sum of molecules. Let $I\subset\{1,\ldots,n\}$ such that $\sum d(x_{i},y_{i})<\delta$. By Lemma 4.2 we have for any $w\in W$
$\sum_{i\in I}\left|{a_{i}}\right|$ $\leqslant\left<{f_{I},G\gamma}\right>\leqslant\left|{\left<{f_{I},Gw}\right>}\right|+\left|{\left<{f_{I},G\gamma-Gw}\right>}\right|$
$\leqslant\varepsilon+\left\|{\gamma-w}\right\|$
where the last inequality follows from the uniform integrability and the fact that $\lambda_{T}(\operatorname{supp}(f_{I}))\leqslant\lambda_{T}(\bigcup_{i\in I}[x_{i},y_{i}])<\delta$. Since this holds for every $w\in W$, we obtain $\sum_{i\in I}\left|{a_{i}}\right|\leqslant\varepsilon+\operatorname{dist}(\gamma,W)$.
(ii) implies clearly $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $C>0$ and for all $w\in W$ and for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$, $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\forall I\subset\{1,\ldots,n\}$,
$\sum_{I}d(x_{i},y_{i})<\delta\implies\sum_{I}|a_{i}|\leq\varepsilon+C\varepsilon.$
Therefore (ii) $\implies$ (iii). ∎
[…]
$m_{xy}=\sum_{i=1}^{k}\dfrac{d(u_{i},u_{i+1})}{d(x,y)}m_{u_{i}u_{i+1}}.$
We also require the following lemma.
Lemma 4.3.
Let $\delta,\varepsilon>0$, $R\geq 1$, and $\gamma:=\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}({M})$ be such that $\sum_{i=1}^{n}\left|{a_{i}}\right|\leq R$ and for every $I\subset\{1,\ldots,n\}$ we have
$\sum_{i\in I}d(x_{i},y_{i})<\delta\implies\sum_{i\in I}|a_{i}|<\varepsilon.$
For every $i\in I$, let $(u_{1}^{i},\ldots,u_{k_{i}+1}^{i})$ be a partition of the couple of points $(x_{i},y_{i})$. Then
(2) $\gamma=\sum_{i=1}^{n}a_{i}\sum_{j=1}^{k_{i}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}m_{u_{j}^{i}u_{j+1}^{i}}$
and for every $\mathcal{I}\subseteq\left\{{(i,j):i\in\left\{{1,\ldots,n}\right\},j\in\left\{{1,\ldots,k_{i}}\right\}}\right\}$ we have
$\sum_{(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})<\frac{\delta\varepsilon}{R}\implies\sum_{(i,j)\in\mathcal{I}}|a_{i}|\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}<3\varepsilon+\frac{4\varepsilon^{2}}{R}.$
Proof.
The new representation (2) of $\gamma$ is clear. Now, fix
$\mathcal{I}\subseteq\left\{{(i,j):i\in\left\{{1,\ldots,n}\right\},j\in\left\{{1,\ldots,k_{i}}\right\}}\right\}$
and assume that
$\sum_{(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})<\frac{\delta\varepsilon}{R}.$
Let $J=\left\{{i\in\left\{{1,\ldots,n}\right\}:\exists j\in\left\{{1,\ldots,k_{i}+1}\right\},(i,j)\in\mathcal{I}}\right\}$ and let us divide $J$ into pairwise disjoint subsets. Let
$J_{0}=\left\{{i\in J:d(x_{i},y_{i})\geq\delta}\right\}$
and for $m\in\mathbb{N}$ let
$J_{m}=\left\{{i\in J:d(x_{i},y_{i})<\delta,\frac{1}{2^{m}}<\sum_{j:(i,j)\in\mathcal{I}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}\leq\frac{1}{2^{m-1}}}\right\}.$
Clearly $J_{0},J_{1},\ldots$ are pairwise disjoint and $J=\bigcup_{m=0}^{\infty}J_{m}$. Now for $i\in J_{0}$ we have
$\sum_{j:(i,j)\in\mathcal{I}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}\leq\frac{1}{\delta}\sum_{j:(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})\leq\frac{1}{\delta}\frac{\delta\varepsilon}{R}=\frac{\varepsilon}{R}.$
Thus
$\sum_{i\in J_{0}}\left|{a_{i}}\right|\sum_{j:(i,j)\in\mathcal{I}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}\leq\frac{\varepsilon}{R}\sum_{i\in J_{0}}\left|{a_{i}}\right|\leq\varepsilon.$
For $m\in\mathbb{N}$ we have
$\sum_{i\in J_{m}}d(x_{i},y_{i})<2^{m}\sum_{i\in J_{m}}\sum_{j:(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i}).$
Note that for every $i\in J_{m}$ we have $d(x_{i},y_{i})<\delta$. Thus we can divide the set $J_{m}$ into $l_{m}\geqslant 1$ pairwise disjoint subsets $A_{1}^{m},\ldots,A_{l_{m}}^{m}$ such that $J_{m}=\bigcup_{j=1}^{l_{m}}A_{j}^{m}$, for every $j\in\left\{{1,\ldots,l_{m}}\right\}$ we have $\sum_{i\in A_{j}^{m}}d(x_{i},y_{i})<\delta$, and for any $j,j^{\prime}\in\left\{{1,\ldots,l_{m}}\right\}$ with $j\neq j^{\prime}$ we have $\sum_{i\in A_{j}^{m}\cup A_{j^{\prime}}^{m}}d(x_{i},y_{i})\geq\delta$. Then
$\frac{\delta}{2}(l_{m}-1)\leq\sum_{i\in J_{m}}d(x_{i},y_{i})\leq 2^{m}\sum_{i\in J_{m}}\sum_{j:(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})$
By assumption $\sum_{i\in A_{j}^{m}}\left|{a_{i}}\right|<\varepsilon$ and thus
$\sum_{i\in J_{m}}\left|{a_{i}}\right|=\sum_{j=1}^{l_{m}}\sum_{i\in A_{j}^{m}}\left|{a_{i}}\right|<l_{m}\varepsilon.$
Now
$\sum_{i\in J_{m}}\left|{a_{i}}\right|\sum_{j:(i,j)\in\mathcal{I}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}\leq\frac{1}{2^{m-1}}\sum_{i\in J_{m}}\left|{a_{i}}\right|<\frac{l_{m}\varepsilon}{2^{m-1}}.$
Therefore:
$\sum_{(i,j)\in\mathcal{I}}|a_{i}|\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}$ $=\sum_{i\in J_{0}}|a_{i}|\sum_{j:(i,j)\in\mathcal{I}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}+\sum_{m=1}^{\infty}\sum_{i\in J_{m}}|a_{i}|\sum_{j:(i,j)\in\mathcal{I}}\frac{d(u_{j}^{i},u_{j+1}^{i})}{d(x_{i},y_{i})}$
$<\varepsilon+\sum_{m=1}^{\infty}\frac{l_{m}\varepsilon}{2^{m-1}}=3\varepsilon+\sum_{m=1}^{\infty}\frac{(l_{m}-1)\varepsilon}{2^{m-1}}$
$\leq 3\varepsilon+\sum_{m=1}^{\infty}\frac{\varepsilon}{2^{m-1}}\frac{2}{\delta}2^{m}\sum_{i\in J_{m}}\sum_{j:(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})$
$=3\varepsilon+\frac{4\varepsilon}{\delta}\sum_{m=1}^{\infty}\sum_{i\in J_{m}}\sum_{j:(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})$
$\leq 3\varepsilon+\frac{4\varepsilon}{\delta}\sum_{(i,j)\in\mathcal{I}}d(u_{j}^{i},u_{j+1}^{i})$
$\leq 3\varepsilon+\frac{4\varepsilon^{2}}{R}.$
Proof of $(iii)\implies(i)$ of Proposition 4.1.
Let $C>0$ be obtained by $(iii)$. Let $\varepsilon>0$ be fixed and $\delta\in(0,1)$ be obtained by(iii)(iii). We set $R=\sup_{w\in W}\|w\|+C\varepsilon$. Consider a measurable subset $B\subset T$ with $\lambda_{T}(B)<\frac{\delta(\varepsilon+C\varepsilon)}{R}$. We will find a bound for $\int_{B}|I(w)|\,d\lambda_{T}$; to do so, by definition of the length measure (see [16]), we may assume that $B$ is a finite union of segments, $B=\bigcup_{k=1}^{N}[a_{k},b_{k}]$, with the sets $(a_{k},b_{k})$ being pairwise disjoint. Pick any $w\in W$ and let $\gamma=\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(\operatorname{supp}(w)\cup\left\{{0}\right\})$ be a convex sum of molecules such that $\|w-\gamma\|\leqslant C\varepsilon$. Notice that $\sum_{i=1}^{n}\left|{a_{i}}\right|=\left\|{\gamma}\right\|\leqslant R$. We may assume that all segments $x_{i},y_{i}]$ are disjoint except for possibly the endpoints, by subdividing each segment if necessary, using the convexity argument outlined earlier, and then grouping and adding equal terms together. Similarly, we assume that the sets $x_{i},y_{i}]$ and $a_{k},b_{k}]$ either coincide or satisfy $\lambda_{T}([x_{i},y_{i}]\cap[a_{k},b_{k}])=0$ (at most finitely many points in the intersection). The new representation of $\gamma$ might contain molecules whose endpoints are not in the set $\operatorname{supp}(w)\cup\left\{{0}\right\}$, so we must enlist the help of Lemma 4.3 to see that for every $I\subseteq\left\{{1,\ldots,n}\right\}$ we have
$(\star)\quad\sum_{I}d(x_{i},y_{i})<\frac{\delta(\varepsilon+C\varepsilon)}{R}\implies\sum_{I}|a_{i}|<3(\varepsilon+C\varepsilon)+\frac{4(\varepsilon+C\varepsilon)^{2}}{R}.$
From here, some direct calculations will yield the desired result:
$\int_{B}|G(w)|\,d\lambda_{T}$ $\leqslant C\varepsilon\lambda_{T}(B)+\sum_{k=1}^{N}\int_{a_{k}}^{b_{k}}|G(\gamma)|\,d\lambda_{T}$
$\leqslant C\varepsilon\lambda_{T}(B)+\sum_{k=1}^{N}\sum_{i=1}^{n}\int_{a_{k}}^{b_{k}}|a_{i}|\dfrac{{\mathbf{1}}_{{[x_{i},y_{i}]}}}{d(x_{i},y_{i})}\,d\lambda_{T}$
$<C\varepsilon+\sum_{k=1}^{N}\sum_{i=1}^{n}\left\{\begin{array}[]{cc}|a_{i}|&\text{ if }[x_{i},y_{i}]=[a_{k},b_{k}]\\ 0&\text{ otherwise}\end{array}\right$.
$\leqslant 3\varepsilon+4C\varepsilon+\frac{4(\varepsilon+C\varepsilon)^{2}}{R}.$
The very last inequality follows from
$\sum_{i\,:\,\exists k,[x_{i},y_{i}]=[a_{k},b_{k}]}d(x_{i},y_{i})\leqslant\sum_{k=1}^{N}d(a_{k},b_{k})\leqslant\frac{\delta(\varepsilon+C\varepsilon)}{R}$
and thus property $(\star)$ provides the last estimate. ∎
Next we aim to give other equivalent conditions for $(ii)$ and $(iii)$. To that end, we will benefit from the following definition.
Definition 4.4.
Let $M$ be a metric space. For $\delta>0$ we define the set
$\mathrm{CSM}_{\delta}\left({M}\right)=\left\{{\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(M):\sum_{i=1}^{n}d(x_{i}y_{i})<\delta,\sum_{i=1}^{n}\left|{a_{i}}\right|=\left\|\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\right\|}\right\}.$
Remark 4.5.
Note that, at least for $\mathbb{R}$-trees, we have
$\mathrm{CSM}_{\delta}\left({T}\right)=\left\{{\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T):\sum_{i=1}^{n}d(x_{i},y_{i})<\delta}\right\}.$
Indeed, let $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(M)$ be such that $\sum_{i=1}^{n}d(x_{i},y_{i})<\delta$. We may assume that all segments $x_{i},y_{i}]$ either coincide or their intersections contain a finite amount of points, by subdividing each segment if necessary and using the convexity argument outlined earlier. Note that in this process the sum $\sum_{i=1}^{n}d(x_{i},y_{i})$ does not change. Next we add together the equal terms. The sum can decrease in this process, but not increase, and thus for the new representation we still have the desired inequality. Furthermore, the new representation is a convex sum of molecules (even an $\ell_{1}$ sum, see the comment at the start of Section 4).
The set $\mathrm{CSM}_{\delta}\left({M}\right)$ will be used in the following proposition.
Proposition 4.6.
Let $T$ be a $\mathbb{R}$-tree. For a bounded $W\subset\mathcal{F}(T)$, the following are equivalent:
  1. $(i)$
    $\forall C>0$, $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$, $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\forall I\subset\{1,\ldots,n\}$, $\sum_{I}d(x_{i},y_{i})<\delta$ implies that $\sum_{I}|a_{i}|\leq\varepsilon+C\varepsilon$.
  2. $(ii)$
    $\exists C>0$, $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$, $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\forall I\subset\{1,\ldots,n\}$, $\sum_{I}d(x_{i},y_{i})<\delta$ implies that $\sum_{I}|a_{i}|\leq\varepsilon+C\varepsilon$.
  3. $(iii)$
    $\forall C>0$, $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(\operatorname{supp}(w)\cup\left\{{0}\right\})$, $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\forall I\subset\{1,\ldots,n\}$, $\sum_{I}d(x_{i},y_{i})<\delta$ implies that $\sum_{I}|a_{i}|\leq\varepsilon+C\varepsilon$.
  4. $(iv)$
    $\exists C>0$, $\forall\varepsilon>0$, $\exists\delta>0$ such that for all $w\in W$ and for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(\operatorname{supp}(w)\cup\left\{{0}\right\})$, $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\forall I\subset\{1,\ldots,n\}$, $\sum_{I}d(x_{i},y_{i})<\delta$ implies that $\sum_{I}|a_{i}|\leq\varepsilon+C\varepsilon$.
  5. $(v)$
    $\forall\varepsilon>0$, $\exists\delta>0$ such that $\operatorname{dist}\left(w,\mathrm{CSM}_{\delta}\left({T}\right)\right)\geq\left\|{w}\right\|-\varepsilon$ for every $w\in W$.
  6. $(vi)$
    $\forall\varepsilon>0$, $\exists\delta>0$ such that $\operatorname{dist}\left(w,\mathrm{CSM}_{\delta}\left({\operatorname{supp}(w)\cup\left\{{0}\right\}}\right)\right)\geq\left\|{w}\right\|-\varepsilon$ for every $w\in W$.
Proof.
The implications $(i)\implies(ii)\implies(iv)$, $(i)\implies(iii)\implies(iv)$, and $(v)\implies(vi)$ are obvious, and the implication $(iv)\implies(i)$ follows from Proposition 4.1.
$(vi)\implies(iv)$: We will prove the contrapositive. Assume that $(iv)$ fails. Fix $C<1$ and obtain $\varepsilon>0$ witnessing the failure of $(iv)$. Fix $\delta>0$ and find $w\in W$, a convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(\operatorname{supp}(w)\cup\left\{{0}\right\})$, and $I\subset\{1,\ldots,n\}$ such that $\|w-\sum_{i}a_{i}m_{x_{i}y_{i}}\|\leqslant C\varepsilon$, $\sum_{I}d(x_{i},y_{i})<\delta$, and $\sum_{I}|a_{i}|>\varepsilon+C\varepsilon$. Then $\sum_{I}a_{i}m_{x_{i}y_{i}}\in\mathrm{CSM}_{\delta}\left({\operatorname{supp}(w)\cup\left\{{0}\right\}}\right)$ and
$\left\|{w-\sum_{I}a_{i}m_{x_{i}y_{i}}}\right\|$ $\leq\left\|{w-\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}}\right\|+\sum_{i\notin I}\left|{a_{i}}\right|$
$\leq C\varepsilon+\sum_{i=1}^{n}\left|{a_{i}}\right|-\sum_{I}\left|{a_{i}}\right|<C\varepsilon+\left\|{\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}}\right\|-(\varepsilon+C\varepsilon)$
$\leq C\varepsilon+\left\|{w}\right\|+C\varepsilon-(\varepsilon+C\varepsilon)=\left\|{w}\right\|-(1-C)\varepsilon.$
Hence we have now shown that for $\varepsilon^{\prime}:=(1-C)\varepsilon$ and for every $\delta>0$ there exists $w\in W$ such that
$\operatorname{dist}\left(w,\mathrm{CSM}_{\delta}\left({\operatorname{supp}(w)\cup\left\{{0}\right\}}\right)\right)<\left\|{w}\right\|-\varepsilon^{\prime},$
i.e. the negation of the condition $(vi)$.
$(i)\implies(v)$: We will prove the contrapositive. Assume that $(v)$ fails and let $\varepsilon>0$ witness its failure. Fix $\delta>0$ and let $w\in W$ be such that $\operatorname{dist}\left(w,\mathrm{CSM}_{\delta}\left({T}\right)\right)<\left\|{w}\right\|-\varepsilon$. Let $\gamma:=\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathrm{CSM}_{\delta}\left({T}\right)$ be such that $\sum_{i=1}^{n}d(x_{i},y_{i})<\delta$ and $\left\|{w-\gamma}\right\|<\left\|{w}\right\|-\varepsilon$. Let $\sum_{i=1}^{k}b_{i}m_{u_{i}v_{i}}\in\mathcal{F}(T)$ be a convex sum of molecules such that
$\left\|{w-\sum_{i=1}^{k}b_{i}m_{u_{i}v_{i}}}\right\|<\frac{\varepsilon}{4}.$
Then $\left\|{\gamma-\sum_{i=1}^{k}b_{i}m_{u_{i}v_{i}}}\right\|<\left\|{w}\right\|-\frac{3\varepsilon}{4}$. By using the convexity argument introduced earlier, we may assume that $u_{i},v_{i}]$ and $x_{j},y_{j}]$ either coincide or are disjoint, except possibly at endpoints, and that all pairs $(u_{i},v_{i})$ are different. Let
$I:=\left\{{i\in\left\{{1,\ldots,k}\right\}:\exists j\in\left\{{1,\ldots,n}\right\},[u_{i},v_{i}]=[x_{j},y_{j}]}\right\}.$
Then $\sum_{I}d(u_{i},v_{i})\leq\sum_{i=1}^{n}d(x_{i},y_{i})<\delta$ and
$\left\|{w}\right\|-\frac{3\varepsilon}{4}>\left\|{\gamma-\sum_{i=1}^{k}b_{i}m_{u_{i}v_{i}}}\right\|\geq\sum_{i\notin I}\left|{b_{i}}\right|=\sum_{i=1}^{k}\left|{b_{i}}\right|-\sum_{I}\left|{b_{i}}\right|>\left\|{w}\right\|-\frac{\varepsilon}{4}-\sum_{I}\left|{b_{i}}\right|,$
where the second inequality follows from the fact that molecules with disjoint support are in $\ell_{1}$ sum, see the start of Section 4. This yields $\sum_{I}\left|{b_{i}}\right|>\frac{\varepsilon}{2}$. As $\delta>0$ was arbitrary, we now get a contradiction with condition $(i)$ when picking $C=1$ and $\varepsilon=\frac{\varepsilon}{4}$. ∎
Several of these implications actually hold for arbitrary metric spaces: $(i)\implies(ii)\implies(iv)$, $(i)\implies(iii)\implies(iv)$, $(v)\implies(vi)\implies(iv)$ and $(v)\implies(ii)$. This leads to the following question.
Question 1.
Do the equivalences in Proposition 4.6 hold for arbitrary metric spaces? Does weak pre-compactness imply any (or all) of them?
We now turn to condition $(b)$ in the definition of equi-integrability mentioned above. It corresponds to being “almost finitely generated” in the following sense.
Proposition 4.7.
Let $T$ be an $\mathbb{R}$-tree. For a bounded $W\subset\mathcal{F}(T)$, the following are equivalent:
  1. $(i)$
    For any $\varepsilon>0$, there exists a measurable set $B\subset T$ with $\lambda_{T}(B)<\infty$ and $\int_{T\setminus B}|Gw|\,d\lambda_{T}<\varepsilon$ for all $w\in W$.
  2. $(ii)$
    For any $\varepsilon>0$, there exists a finitely generated subtree $T^{\prime}:=\bigcup_{i=1}^{n}[0,z_{i}]\subset T$ such that $W\subset\mathcal{F}(T^{\prime})+\varepsilon B_{\mathcal{F}(T)}$.
Proof.
$(i)\implies(ii)$: Fix $\varepsilon>0$. Let $B$ be a measurable subset as in $(i)$. By the definition of $\lambda_{T}$, we may assume that $B$ is a finite union of disjoint segments: $B=\bigcup_{i=1}^{N}S_{i}$. Let $T^{\prime}$ be the subtree of $T$ generated by 0 and the segments $(S_{i})_{i=1}^{N}$. The resulting tree $T^{\prime}$ can be expressed as $T^{\prime}=\bigcup_{i=1}^{n}[0,z_{i}]$.
Now, consider $w\in W$. By the assumption, we have:
$\int_{T}|Gw|\,d\lambda_{T}-\int_{T}{\mathbf{1}}_{{B}}|Gw|\,d\lambda_{T}=\int_{T\setminus B}|Gw|\,d\lambda_{T}<\varepsilon.$
Thus, if we define $w^{\prime}:=G^{-1}({\mathbf{1}}_{{B}}Gw)$, then we have that $w^{\prime}\in\mathcal{F}(T^{\prime})$ and:
$\|w-w^{\prime}\|$ $=\|Gw-Gw^{\prime}\|=\int_{T}|Gw-Gw^{\prime}|\,d\lambda_{T}$
$=\int_{T}|Gw-{\mathbf{1}}_{{B}}Gw|\,d\lambda_{T}=\int_{T\setminus B}|Gw|\,d\lambda_{T}<\varepsilon.$
$(ii)\implies(i)$: Fix $\varepsilon>0$. Let $T^{\prime}$ be a finitely generated subtree as in $(ii)$. Setting $B=T^{\prime}$, we have a measurable subset of $T$ with $\lambda_{T}(B)<\infty$. Consider $w\in W$. By the assumption, there exists $w^{\prime}\in\mathcal{F}(B)$ such that $\|w-w^{\prime}\|\leqslant\varepsilon$. The following calculation completes the proof:
$\int_{T\setminus B}|Gw|\,d\lambda_{T}=\int_{T\setminus B}|Gw-Gw^{\prime}|\,d\lambda_{T}\leqslant\int_{T}|Gw-Gw^{\prime}|\,d\lambda_{T}=\|w-w^{\prime}\|\leqslant\varepsilon.$
Combining Proposition 4.1 with Proposition 4.7 yields the next characterization (Theorem D).
Corollary 4.8.
Let $T$ be an $\mathbb{R}$-tree. If $W\subset\mathcal{F}(T)$ then $W$ is relatively weakly compact if and only if it is bounded and satisfies:
  1. $(a)$
    $\forall\varepsilon>0$, $\exists\delta>0$ such that for every convex sum of molecules $\sum_{i=1}^{n}a_{i}m_{x_{i}y_{i}}\in\mathcal{F}(T)$, $\forall I\subset\{1,\ldots,n\}$,
    $\sum_{i\in I}d(x_{i},y_{i})<\delta\implies\sum_{i\in I}\left|{a_{i}}\right|\leqslant\varepsilon+\operatorname{dist}(\gamma,W).$
  2. $(b)$
    For any $\varepsilon>0$, there exists a finitely generated subtree $T^{\prime}:=\bigcup_{i=1}^{n}[0,z_{i}]\subset T$ such that $W\subset\mathcal{F}(T^{\prime})+\varepsilon B_{\mathcal{F}(T)}$.
Remark 4.9.
The condition (a) can be replaced with any of the equivalent conditions from Proposition 4.6.
Remark 4.10.
While condition (a) in Corollary 4.8 is meaningful in any metric space, condition (b) is intrinsically tied to the tree structure of the underlying space. From the perspective of potential generalizations, it would be valuable to identify an alternative condition that does not explicitly rely on the tree structure of $T$. In particular, it remains unclear whether condition (b) can be substituted by uniform regularity.
Question 2.
Can condition $(b)$ in Corollary 4.8 be replaced with uniform regularity?
The next example shows that (b) cannot always be replaced by tightness.
Example 4.11.
Let $M=\{0\}\cup\bigcup_{n=1}^{\infty}\{x^{n}_{1},\ldots,x^{n}_{n}\}$ with a tree metric $d$ such that 0 is unique branching point and each non-zero point is sitting on its own branch with $d(x^{n}_{i},0)=\frac{1}{n}$. Let $\mu_{n}=\frac{1}{n}(m_{x^{n}_{1}0}+\ldots+m_{x^{n}_{n}0})=\delta(x^{n}_{1})+\ldots+\delta(x^{n}_{n})$. Let $T$ be the smallest $\mathbb{R}$-tree which contains $M$. Then $W:=\left\{{\mu_{n}:n\in M}\right\}$ is isometrically equivalent to the $\ell_{1}$-basis, and therefore $W$ is not weakly compact. Notice also that $W$ is tight since it is bounded in $\mathcal{F}(M)$ and $M$ is compact. Next, we will show that condition $(vi)$ from Proposition 4.6, and thus condition $(a)$ from Corollary 4.8, is satisfied. To that end, fix $\varepsilon>0$ and choose $\delta=\varepsilon$. Fix $\mu_{n}\in W$ and a convex sum of molecules $\gamma:=\sum_{i=1}^{k}a_{i}m_{y_{i}z_{i}}\in\mathrm{CSM}_{\delta}\left({\operatorname{supp}(\mu_{n})\cup\left\{{0}\right\}}\right)$. Without loss of generality we may assume that for every $i\in\left\{{1,\ldots,k}\right\}$ either $y_{i}=0$ or $z_{i}=0$, and that each $x_{i}^{n}$ appears once at most. Let $I:=\left\{{i\in\left\{{1,\ldots,n}\right\}:x_{i}^{n}\in\left\{{y_{1},z_{1},\ldots,y_{k},z_{k}}\right\}}\right\}$. Then
$\frac{\left|{I}\right|}{n}=\sum_{I}d(y_{i},z_{i})\leq\sum_{i=1}^{k}d(y_{i},z_{i})<\delta$
and
$\left\|{\mu_{n}-\sum_{i=1}^{k}a_{i}m_{y_{i}z_{i}}}\right\|\geq\sum_{i\notin I}\frac{1}{n}=1-\frac{\left|{I}\right|}{n}>1-\delta=1-\varepsilon.$
This shows that tightness and condition $(a)$ are not sufficient for weak compactness. However, in this specific example, condition $(b)$ in Corollary 4.8 can be replaced by uniform regularity. Indeed, let $W\subset\mathcal{F}(T)$ and $\varepsilon>0$. Applying uniform regularity to the open set $U=M\setminus\{0\}$, we obtain a compact set $K\subset U$ such that
$W\subset\mathcal{F}(K\cup\{0\})+\varepsilon B_{\mathcal{F}(M)}.$
Since $K\subset T$ is compact, there exists a finitely generated subtree $T^{\prime}$ such that $K\cup\{0\}\subset T^{\prime}$. Hence, uniform regularity implies condition $(b)$. Finally, if $W$ satisfies conditions $(a)$ and $(b)$, then it is relatively weakly compact, and therefore uniformly regular by Corollary 3.1. The same argument goes through for trees with finitely many branching points $\mathrm{Br}(T)$. In that case take $U=M\setminus\mathrm{Br}(T)$ and proceed as above.
5. Sequences of sums of “small molecules” contain an $\ell_{1}$-basis
The previous section highlights the significance of approximation by convex series of molecules. Informally, within a weakly compact set, the weight of molecules formed by pairs of points in close proximity should be controlled in a uniform way. In this section, we establish that this condition is, in a sense, necessary. Specifically, consider a sequence of (finite or countable) linear combinations of molecules, denoted by $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}$, satisfying $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ as $n\to\infty$ and $\sup_{n}\sum_{k}|\lambda_{k}^{n}|<\infty$. We will demonstrate that this sequence contains an $\ell_{1}$-basis, thereby proving that such a sequence cannot be contained in a weakly precompact set.
We shall start by introducing a definition of $\delta$-cyclical monotonicity between subsets of $\widetilde{M}:=\left\{{(x,y):x,y\in M,x\neq y}\right\}$. For a subset of $\widetilde{M}$, the concept of cyclical monotonicity can be found e.g. in [11]: the set $A\subset\widetilde{M}$ is cyclically monotonic if, for every finite sequence of pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A$, we have
$\sum_{i=1}^{n}d(x_{i},y_{i+1})\geqslant\sum_{i=1}^{n}d(x_{i},y_{i}),$
where $y_{n+1}=y_{1}$. This concept was generalized in [18] by introducing a parameter $\delta\in(0,1]$: a subset $A\subset\widetilde{M}$ is $\delta$-cyclically monotonic, if for any finite sequence of pairs $(x_{1},y_{1})$, $\ldots$ , $(x_{n},y_{n})\in A$, we have
$\sum_{i=1}^{n}\min\left\{{d(x_{i},y_{i+1})-\delta d(x_{i},y_{i}),d(y_{i},y_{i+1})}\right\}\geq 0,$
where $y_{n+1}:=y_{1}$. Note that $A$ is cyclically monotonic if and only if it is 1-cyclically monotonic. The motivation for both [11] and [18] to consider these definitions was to find a lower estimate for the norm of the sum of molecules. In particular, by [18, Proposition 2.2] (or by [11, Theorem 2.4] for the case $\delta=1$), a set $A\subset\widetilde{M}$ is $\delta$-cyclically monotonic if and only if there exists $f\in S_{\operatorname{Lip}_{0}(M)}$ such that $\langle f,m_{xy}\rangle\geq\delta$ for every $(x,y)\in A$.
Definition 5.1.
Let $A,B\subset\widetilde{M}$ and let $\delta\in(0,1]$. We say that the sets $A$ and $B$ are $\delta$-cross-cyclically monotonic, if for any finite sequence of pairwise distinct pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A\cup B$, which contains at least one pair from each of $A$ and $B$, we have
(3) $\sum_{i=1}^{n}\min\left\{{d(x_{i},y_{i+1})-\delta d(x_{i},y_{i}),d(y_{i},y_{i+1})}\right\}\geq 0,$
where $y_{n+1}:=y_{1}$. When the sets $A$ and $B$ are 1-cross-cyclically monotonic, the inequality (3) reduces to
$\sum_{i=1}^{n}\left(d(x_{i},y_{i+1})-d(x_{i},y_{i})\right)\geq 0,$
and in this case, we simply say that the sets $A$ and $B$ are cross-cyclically monotonic.
Remark 5.2.
Notice that when $A$ and $B$ are $\delta$-cross-cyclically monotonic (with $\delta>0$), then $(x,y)\in A$ implies $(y,x)\notin B$. Nevertheless this does not prevent both $(x,y)$ and $(y,x)$ from belonging to $A$. Indeed, let $M=\left\{{x,y,a,b}\right\}$ such that $d(x,y)=d(a,b)=1$ and every other non-zero distance is $2$. Then $A=\left\{{(x,y),(y,x)}\right\}$ and $B=\left\{{(a,b),(b,a)}\right\}$ are cross-cyclically monotonic, since for any sequence of pairwise distinct elements $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A\cup B$, $n=2,3,4$, that contains elements from both sets, we have
$\sum_{i=1}^{n}d(x_{i},y_{i+1})\geq 4\geq\sum_{i=1}^{n}d(x_{i},y_{i}),$
where $y_{i+1}=y_{1}$. Note that $A$ and $B$ are clearly not $\delta$-cyclically monotonic for any $\delta>0$. In particular, cross-cyclical monotonicity places no cyclical-monotonicity constraint on the individual components.
Remark 5.3.
We need the assumption that the pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are pairwise distinct to avoid the situation when we select the same pair twice from the same set. However, it is straightforward to check that, when the sets $A$ and $B$ are cross-cyclically monotonic, then the inequality (3) holds also for all such sequences $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A\cup B$, where whenever two pairs $(x_{k},y_{k})$ and $(x_{k^{\prime}},y_{k^{\prime}})$ are equal, then they belong to different sets $A$ and $B$.
We now extend [11, Theorem 2.4] and [18, Proposition 2.2] to the setting of mutually $\delta$-cyclically monotonic sets.
Proposition 5.4.
Let $\delta\in(0,1]$ and let $A_{1},A_{2}\subset\widetilde{M}$ be $\delta$-cross-cyclically monotonic. Let $\lambda^{i}\in\ell_{1}^{+}(A_{i})$, and denote $\gamma_{i}=\sum_{(x,y)\in A_{i}}\lambda^{i}_{(x,y)}m_{xy}$, for $i=1,2$. Then there exists $h\in S_{\operatorname{Lip}_{0}(M)}$ such that $\left<{h,\gamma_{i}}\right>\geqslant\delta\left\|{\gamma_{i}}\right\|$ for $i=1,2$. In particular $\left\|{\gamma_{1}+\gamma_{2}}\right\|\geqslant\delta(\left\|{\gamma_{1}}\right\|+\left\|{\gamma_{2}}\right\|)$.
[…]
Lemma 5.5.
Let $I$ be a set and $\beta_{ij}$, $i,j\in I$, be real numbers such that $\beta_{ii}=0$ for every $i\in I$. Then the following statements are equivalent:
  • (i)
    There exist numbers $\alpha_{i}$, $i\in I$ such that
    $\alpha_{i}\leqslant\alpha_{j}+\beta_{ij}$
    for $i,j\in I$.
  • (ii)
    For every finite sequence $i_{1},\ldots,i_{m}$ in $I$ we have
    $\beta_{i_{1}i_{2}}+\beta_{i_{2}i_{3}}\ldots+\beta_{i_{m}i_{1}}\geqslant 0,$
  • (iii)
    For every injective finite sequence $i_{1},\ldots,i_{m}$ in $I$ we have
    $\beta_{i_{1}i_{2}}+\beta_{i_{2}i_{3}}\ldots+\beta_{i_{m}i_{1}}\geqslant 0,$
Proof of Proposition 5.4.
The proof follows similar ideas as [11, Theorem 2.4] and [18, Proposition 2.2]. Let $f,g\in S_{\operatorname{Lip}_{0}(M)}$ be such that $\langle f,\gamma_{1}\rangle=\left\|{\gamma_{1}}\right\|$ and $\langle g,\gamma_{2}\rangle=\left\|{\gamma_{2}}\right\|$. Write $A_{1}=\left\{{(x_{i},y_{i}):i\in I_{1}}\right\}$ and $A_{2}=\left\{{(x_{i},y_{i}):i\in I_{2}}\right\}$, where $I_{1}$ and $I_{2}$ are disjoint sets of indices. Let $I=I_{1}\cup I_{2}$. For $i,j\in I$, let
$\beta_{ij}:=\begin{cases}\min\left\{{d(x_{i},y_{j})-\delta\left(f(x_{i})-f(y_{i})\right),d(y_{i},y_{j})}\right\},&\text{ if }i\in I_{1}\\ \min\left\{{d(x_{i},y_{j})-\delta\left(g(x_{i})-g(y_{i})\right),d(y_{i},y_{j})}\right\},&\text{ if }i\in I_{2}.\end{cases}$
Clearly $\beta_{jj}=0$ for all $j\in I$. Let us consider pairwise distinct indices $i_{1},\ldots,i_{n}\in I$. Our aim is to show that
$\sum_{k=1}^{n}\beta_{i_{k}i_{k+1}}\geq 0,$
where we denote $i_{n+1}:=i_{1}$. First consider the case when $i_{1},\ldots,i_{n}\in I_{1}$. Then
$\sum_{k=1}^{n}\beta_{i_{k}i_{k+1}}$ $=\sum_{k=1}^{n}\min\left\{{d(x_{i_{k}},y_{i_{k+1}})-\delta\left(f(x_{i_{k}})-f(y_{i_{k}})\right),d(y_{i_{k}},y_{i_{k+1}})}\right\}$
$\geq\delta\sum_{k=1}^{n}\min\left\{{d(x_{i_{k}},y_{i_{k+1}})-f(x_{i_{k}})+f(y_{i_{k}}),d(y_{i_{k}},y_{i_{k+1}})}\right\}$
$\geq\delta\sum_{k=1}^{n}\left(-f(y_{i_{k+1}})+f(y_{i_{k}})\right)=0.$
The case when $i_{1},\ldots,i_{n}\in I_{2}$ is analogous. Now assume that $i_{1},\ldots,i_{n}$ contain indices from both sets $I_{1}$ and $I_{2}$. For all $i,j\in I$, we have
$\beta_{ij}\geq\min\left\{{d(x_{i},y_{j})-\delta d(x_{i},y_{i}),d(y_{i},y_{j})}\right\},$
and thus we obtain $\sum_{k=1}^{n}\beta_{i_{k}i_{k+1}}\geq 0$ from the fact that the sets $A_{1}$ and $A_{2}$ are $\delta$-cross-cyclically monotonic and from Remark 5.3. Indeed, notice that if for $i,j\in I$ with $i\neq j$ we have $(x_{i},y_{i})=(x_{j},y_{j})$ then necessarily $(x_{i},y_{i})\in A_{1}$ and $(x_{j},y_{j})\in A_{2}$ or vice versa. Next, from Lemma 5.5 we obtain real numbers $\alpha_{i}$, $i\in I$ such that
$\alpha_{i}\leq\alpha_{j}+\beta_{ij}$
for all $i,j\in I$. Define $h:M\rightarrow\mathbb{R}$ by
$h(u)=\inf_{i\in I}(\alpha_{i}+d(u,y_{i})).$
As the infimum of 1-Lipschitz functions, $h$ is 1-Lipschitz. The properties of $\alpha_{i}$, $i\in I$, are not affected by adding a constant, thus we may additionally assume they were chosen in such a way that $h(0)=0$ and thus $h\in B_{\operatorname{Lip}_{0}(M)}$. It remains to show that $\langle h,\gamma_{i}\rangle\geq\delta\left\|{\gamma_{i}}\right\|$, for $i=1,2$. Fix $i\in I_{1}$ and $\varepsilon>0$. There exists $j\in I$ such that $h(x_{i})\geq\alpha_{j}+d(x_{i},y_{j})-\varepsilon$. Furthermore, $h(y_{i})\leq\alpha_{i}$, and thus
$h(x_{i})-h(y_{i})\geq\alpha_{j}+d(x_{i},y_{j})-\varepsilon-\alpha_{i}\geq d(x_{i},y_{j})-\beta_{ij}-\varepsilon\geq\delta\left(f(x_{i})-f(y_{i})\right)-\varepsilon.$
As $\varepsilon>0$ was arbitrary, we obtain $h(x_{i})-h(y_{i})\geq\delta\left(f(x_{i})-f(y_{i})\right)$ for every $i\in I_{1}$, and thus $\langle h,\gamma_{1}\rangle\geq\delta\langle f,\gamma_{1}\rangle=\delta\left\|{\gamma_{1}}\right\|$. By the same argument, we obtain $h(x_{i})-h(y_{i})\geq\delta\left(g(x_{i})-g(y_{i})\right)$ for every $i\in I_{2}$ and thus $\langle h,\gamma_{2}\rangle\geq\delta\langle g,\gamma_{2}\rangle=\delta\left\|{\gamma_{2}}\right\|$. This concludes the proof. ∎
Lemma 5.6.
Let $M$ be a pointed metric space. Let $\mu:=\sum_{k}\lambda_{k}m_{x_{k}y_{k}}\in\mathcal{F}(M)$ where $(\lambda_{k})\in\ell_{1}$ and let $\gamma\in\mathcal{F}(M)$. Let $\delta:=\sum_{k}d(x_{k},y_{k})$ and
$C:=\inf\{d(x,y):\,x\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\},\,y\in\left\{{x_{1},y_{1},x_{2},y_{2},\ldots}\right\}\},$
If $C>\delta$, then $\left\|{\gamma+\mu}\right\|\geq\frac{C}{C+\delta}(\|\gamma\|+\|\mu\|)$.
Proof.
By an obvious approximation argument, we may assume that $\gamma$ and $\mu$ are finitely supported, with $\mu$ written as $\mu:=\sum_{k=1}^{n}\lambda_{k}m_{x_{k}y_{k}}$. We may additionally assume $\lambda_{k}>0$ by interchanging $x_{k}$ and $y_{k}$ whenever necessary. Let $A=\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ and $B=\left\{{x_{1},y_{1},\ldots,x_{n},y_{n}}\right\}$. We may further assume that $\gamma$ is written as a convex sum of molecules from points in $A$, say $m_{x_{n+1}y_{n+1}},\ldots,m_{x_{N}y_{N}}$, where $x_{n+1},y_{n+1},\ldots,x_{N},y_{N}\in A$. Finally, we assume without loss of generality that $M=A\cup B$. Note that $A\cap B=\emptyset$ as $d(A,B)=C>\delta>0$.
Define a new metric $b$ on $M$ by
$b(x,y)=\begin{cases}d(x,y)&\text{ if $x,y\in A$ or $x,y\in B$}\\ d(x,y)+\delta&\text{ otherwise.}\end{cases}$
To prove that it is indeed a metric, it is enough to check the triangle inequality. Let $x,y,z\in M$ be different. If $x,z$ are in the same set $A$ or $B$ then
$b(x,z)=d(x,z)\leqslant d(x,y)+d(y,z)\leqslant b(x,y)+b(y,z).$
Otherwise, $y$ must be in the same set as exactly one of them, say $x$, then
$b(x,z)=d(x,z)+\delta\leqslant d(x,y)+d(y,z)+\delta=b(x,y)+b(y,z).$
So $b$ is indeed a metric on $M$ and, moreover, it is equivalent to $d$: if $d(x,y)\neq b(x,y)$ then $d(x,y)\geqslant C$, thus
$b(x,y)=d(x,y)+\delta\leqslant\left({1+\frac{\delta}{C}}\right)d(x,y).$
Therefore $d\leqslant b\leqslant(1+\delta/C)d$. In particular $\mathcal{F}({M,d})$ and $\mathcal{F}({M,b})$ are $(1+\delta/C)$-isomorphic.
Claim: The sets $\left\{{(x_{1},y_{1}),\ldots,(x_{n},y_{n})}\right\}$ and $\left\{{(x_{n+1},y_{n+1}),\ldots,(x_{N},y_{N})}\right\}$ are cross-cyclically monotonic in the metric $b$.
Let us assume that the claim is proved. By Proposition 5.4 we get
$\left\|{\mu+\gamma}\right\|_{\mathcal{F}(M,b)}=\left\|{\mu}\right\|_{\mathcal{F}(M,b)}+\left\|{\gamma}\right\|_{\mathcal{F}(M,b)}.$
Furthermore, by definition of $b$, we have $\left\|{\mu}\right\|_{\mathcal{F}(M,b)}=\left\|{\mu}\right\|_{\mathcal{F}(M,d)}$ and $\left\|{\gamma}\right\|_{\mathcal{F}(M,b)}=\left\|{\gamma}\right\|_{\mathcal{F}(M,d)}$, and thus
$\|\mu\|_{\mathcal{F}(M,d)}+\|\gamma\|_{\mathcal{F}(M,d)}\leq\left\|{\mu+\gamma}\right\|_{\mathcal{F}(M,b)}\leq\left({1+\frac{\delta}{C}}\right)\left\|{\mu+\gamma}\right\|_{\mathcal{F}(M,d)}.$
To conclude the proof, it only remains to prove the claim. Note that, since $\gamma$ is a convex sum of the molecules $m_{x_{n+1}y_{n+1}},\dots,m_{x_{N}y_{N}}$, by [11, Theorem 2.4] the sequence $(x_{k},y_{k})_{k=n+1}^{N}$ is cyclically monotonic with respect to $d$, hence also with respect to $b$. Now let us consider $i_{1},\ldots,i_{m}\in\{1,\ldots,N\}$, with $m\geqslant 2$ and indices from both sets $\{1,\ldots,n\}$ and $\{n+1,\ldots,N\}$. Note that we may assume that the indices $i_{1},\ldots,i_{m}$ are pairwise distinct. Let $j_{1}<\ldots<j_{l}$ be such that the indices $i_{j_{k}}$, $1\leqslant k\leqslant l$, are those which belong to the set $\left\{{n+1,\ldots,N}\right\}$, and let us write $I=\{j_{1},\ldots,j_{l}\}$. To simplify the notation in the next estimates, we let $j_{l+1}:=j_{1}$, $i_{m+1}:=i_{1}$. Using the cyclic monotonicity of $(x_{k},y_{k})_{k=n+1}^{N}$ and the definition of $\delta$, we first obtain:
$\sum_{k=1}^{m}b(x_{i_{k}},y_{i_{k}})$ $=\sum_{k=1}^{l}d(x_{i_{j_{k}}},y_{i_{j_{k}}})+\sum_{\begin{subarray}{c}k=1\\ k\not\in I\end{subarray}}^{m}d(x_{i_{k}},y_{i_{k}})$
$\leqslant\sum_{k=1}^{l}d(x_{i_{j_{k}}},y_{i_{j_{k+1}}})+\delta$
Next, using the triangle inequality, we get:
$\sum_{k=1}^{m}b(x_{i_{k}},y_{i_{k}})$ $\leqslant\sum_{k=1}^{l}\left(\sum_{k^{\prime}=j_{k}}^{j_{k+1}-1}d(x_{i_{k^{\prime}}},y_{i_{k^{\prime}+1}})+\sum_{k^{\prime}=j_{k}+1}^{j_{k+1}-1}d(x_{i_{k^{\prime}}},y_{i_{k^{\prime}}})\right)+\delta$
$\leqslant\sum_{k=1}^{m}d(x_{i_{k}},y_{i_{k+1}})+\sum_{k=1}^{n}d(x_{k},y_{k})+\delta$
$\leqslant\sum_{k=1}^{m}d(x_{i_{k}},y_{i_{k+1}})+2\delta\leqslant\sum_{k=1}^{m}b(x_{i_{k}},y_{i_{k+1}}).$
This ends the proof of the claim, and the lemma. ∎
Lemma 5.7.
Let $M$ be a pointed metric space. Let $\mu:=\sum_{k}\lambda_{k}m_{x_{k}y_{k}}\in\mathcal{F}(M)$ where $(\lambda_{k})\in\ell_{1}$ and let $\gamma\in\mathcal{F}(M)$. Let $r>0$ be such that for every $k$ there exists $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ such that
$x_{k},y_{k}\in B(u,r).$
Then $\left\|{\gamma+\mu}\right\|\geq\frac{D-2r}{D+2r}(\|\gamma\|+\|\mu\|)$, where
$D=\inf\left\{{d(x,y):x\neq y\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}}\right\}.$
Proof.
As in the previous lemma, we may assume that $\gamma$ and $\mu$ are finitely supported, and that $\mu:=\sum_{k=1}^{n}\lambda_{k}m_{x_{k}y_{k}}$. The case when $r\geq\frac{D}{2}$ is trivial, thus we assume $r<\frac{D}{2}$. Without loss of generality, we may assume that $M=\operatorname{supp}(\gamma)\cup\left\{{0,x_{1},y_{1},\ldots,x_{n},y_{n}}\right\}$. Then the sets $B(u,r)$, $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ cover the entire $M$. Let $f,g\in S_{\operatorname{Lip}_{0}(M)}$ be such that $\langle f,\mu\rangle=\left\|{\mu}\right\|$ and $\langle g,\gamma\rangle=\left\|{\gamma}\right\|$. Define $h:M\rightarrow\mathbb{R}$ by
$h\mathord{\upharpoonright}_{B(u,r)}=g(u)-f(u)+f\mathord{\upharpoonright}_{B(u,r)}$
for every $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$. Since $r<\frac{D}{2}$, the sets $B(u,r)$, $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ are pairwise disjoint and thus $h$ is well defined, and $h(0)=0$. Next we estimate the Lipschitz constant of $h$. Fix $v,v^{\prime}\in M$ and let $u,u^{\prime}\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ be such that $v\in B(u,r)$ and $v^{\prime}\in B(u^{\prime},r)$. If $u=u^{\prime}$, then
$\left|{h(v)-h(v^{\prime})}\right|=\left|{f(v)-f(v^{\prime})}\right|\leq d(v,v^{\prime}).$
Otherwise $u\neq u^{\prime}$ and thus by triangle inequality
$d(v,v^{\prime})\geq d(u,u^{\prime})-d(u,v)-d(u^{\prime},v^{\prime})\geq D-2r,$
giving us
$\left|{h(v)-h(v^{\prime})}\right|$ $=\left|{g(u)-g(u^{\prime})+f(v)-f(u)+f(u^{\prime})-f(v^{\prime})}\right|$
$\leq d(u,u^{\prime})+d(v,u)+d(u^{\prime},v^{\prime})$
$\leq D+2r\leq\frac{D+2r}{D-2r}d(v,v^{\prime}).$
Therefore $h\in\operatorname{Lip}_{0}(M)$ with $\left\|{h}\right\|\leq\frac{D+2r}{D-2r}$. On the other hand $h(u)=g(u)$ for $u\in\operatorname{supp}(\gamma)$ and thus $\langle h,\gamma\rangle=\langle g,\gamma\rangle=\left\|{\gamma}\right\|$. Furthermore, for any $k\in\left\{{1,\ldots,n}\right\}$ there exists $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ such that $x_{k},y_{k}\in B(u,r)$ and thus
$h(x_{k})-h(y_{k})=f(x_{k})-f(y_{k})$
giving us $\langle h,\mu\rangle=\langle f,\mu\rangle=\left\|{\mu}\right\|$. Hence
$\left\|{\gamma+\mu}\right\|\geq\frac{1}{\left\|{h}\right\|}\langle h,\gamma+\mu\rangle\geq\frac{D-2r}{D+2r}(\|\gamma\|+\|\mu\|).$
Using the two previous lemmas, we obtain the following generalization of [20, Theorem 2.6], which covers the case where each $\gamma_{n}$ is a molecule.
Theorem 5.8.
Let $(\gamma_{n})$ be a sequence in $S_{\mathcal{F}({M})}$ such that each $\gamma_{n}$ is expressed as follows:
$\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}.$
If $\lim_{n\to\infty}\sum_{k}d(x_{k}^{n},y_{k}^{n})=0,$ and $\sup_{n}\sum_{k}\left|{\lambda_{k}^{n}}\right|<\infty$, then $\lim_{n}\left\|{\gamma+\gamma_{n}}\right\|=\left\|{\gamma}\right\|+1$ for all $\gamma\in\mathcal{F}({M})$.
Proof.
By an easy approximation argument, we may assume without loss of generality that each $\gamma_{n}$ is finitely supported, and we write:
$\gamma_{n}=\sum_{k=1}^{N_{n}}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}.$
Further, we may assume $\lambda_{k}^{n}>0$ by interchanging $x_{k}^{n}$ and $y_{k}^{n}$ whenever necessary. We may also assume that $\gamma$ is finitely supported. Fix $\varepsilon\in(0,1/3)$. Let
$D:=\min\left\{{d(x,y):x\neq y\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}}\right\}.$
Let $m\in\mathbb{N}$ be such that $\frac{1}{m}\sup_{n}\sum_{k}\left|{\lambda_{k}^{n}}\right|<\varepsilon$ and let $\delta>0$ be such that $\delta\left(1+\frac{1}{\varepsilon}\right)^{m}+\delta<\varepsilon D$. Next we pick $n\in\mathbb{N}$ large enough so that
$\sum_{k=1}^{N_{n}}d(x_{k}^{n},y_{k}^{n})<\delta.$
Let $A_{0}=\emptyset$,
$A_{i}=\bigcup_{u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}}B\left(u,\delta\left(1+\frac{1}{\varepsilon}\right)^{i}\right)$
for $i=1,\ldots,m$, and $B_{i}=A_{i}\setminus A_{i-1}$ for $i=1,\ldots,m$. Note that the sets $B_{i}$ are pairwise disjoint. Hence each $x_{k}^{n}$ can belong to at most one of these sets. Thus, thanks to our choice of $m$, there must exist $i\in\left\{{1,\ldots,m}\right\}$ such that $\sum_{k\,:\,x_{k}^{n}\in B_{i}}\lambda_{k}^{n}<\varepsilon$. Let
$I$ $=\left\{{k\in\left\{{1,\ldots,N_{n}}\right\}:x_{k}^{n}\notin A_{i}}\right\}$
$J$ $=\left\{{k\in\left\{{1,\ldots,N_{n}}\right\}:x_{k}^{n}\in A_{i-1}}\right\}.$
Then
$\left\|{\sum_{k\in I\cup J}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}}\right\|\geq 1-\sum_{k\notin I\cup J}\lambda_{k}^{n}=1-\sum_{k\,:\,x_{k}^{n}\in B_{i}}\lambda_{k}^{n}>1-\varepsilon.$
Let $\mu=\sum_{k\in J}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}$. For every $k\in J$ there exists $u\in\operatorname{supp}(\gamma\cup\left\{{0}\right\})$ such that
$x_{k}^{n}\in B\left(u,\delta\left(1+\frac{1}{\varepsilon}\right)^{i-1}\right)\subset B\left(u,\varepsilon D-\delta\right)$
and since $d(x_{k}^{n},y_{k}^{n})<\delta$, we get
$x_{k}^{n},y_{k}^{n}\in B(u,\varepsilon D).$
We apply Lemma 5.7 to elements $\mu$ and $\gamma$ and obtain
$\left\|{\gamma+\mu}\right\|\geq\frac{D-2\varepsilon D}{D+2\varepsilon D}\left(\|\gamma\|+\left\|{\mu}\right\|\right)=\frac{1-2\varepsilon}{1+2\varepsilon}\left(\|\gamma\|+\left\|{\mu}\right\|\right).$
Now consider $C:=\frac{\delta}{\varepsilon}-2\delta>\delta$. Since for every $x\in\operatorname{supp}(\gamma+\mu)\cup\left\{{0}\right\}$ there exists $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ such that $x\in B\left(u,\delta\left(1+\frac{1}{\varepsilon}\right)^{i-1}+\delta\right)$ and for every $y\in\left\{{x_{k}^{n},y_{k}^{n}:k\in I}\right\}$ we have $y\notin B\left(u,\delta\left(1+\frac{1}{\varepsilon}\right)^{i}-\delta\right)$, we deduce that
$d(x,y)$ $\geq d(u,y)-d(u,x)\geq\delta\left(1+\frac{1}{\varepsilon}\right)^{i}-\delta-\left(\delta\left(1+\frac{1}{\varepsilon}\right)^{i-1}+\delta\right)$
$=\delta\left(1+\frac{1}{\varepsilon}\right)^{i-1}\frac{1}{\varepsilon}-2\delta\geq\frac{\delta}{\varepsilon}-2\delta=C.$
Thus from Lemma 5.6 we get
$\left\|{\gamma+\mu+\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}}\right\|$ $\geq\frac{C}{C+\delta}\Big(\|\gamma+\mu\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big)$
$=\frac{1-2\varepsilon}{1-\varepsilon}\Big(\|\gamma+\mu\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big).$
Therefore
$\left\|{\gamma+\gamma_{n}}\right\|$ $\geq\left\|{\gamma+\mu+\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}}\right\|-\left\|{\sum_{k\notin I\cup J}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}}\right\|$
$\geq\frac{1-2\varepsilon}{1-\varepsilon}\Big(\|\gamma+\mu\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big)-\varepsilon$
$\geq\frac{(1-2\varepsilon)^{2}}{(1-\varepsilon)(1+2\varepsilon)}\Big(\|\gamma\|+\left\|{\mu}\right\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big)-\varepsilon$
$\geq\frac{(1-2\varepsilon)^{2}}{(1-\varepsilon)(1+2\varepsilon)}\Big(\|\gamma\|+1-\varepsilon\Big)-\varepsilon.$
As $\varepsilon$ was arbitrary, we get $\lim_{n}\left\|{\gamma+\gamma_{n}}\right\|=\left\|{\gamma}\right\|+1$. ∎
This is enough to deduce the already announced main result of this section, Theorem E. We recall that a sequence $(e_{n})\subset S_{X}$ of distinct points in the unit sphere of a Banach space $X$ is called an asymptotically isometric $\ell_{1}$-basis if for every $C\in(0,1)$ there exists $k_{0}\in\mathbb{N}$ such that for every $(a_{k})\in c_{00}$ we have $\left\|{\sum_{k=k_{0}}^{\infty}a_{k}e_{k}}\right\|\geqslant C\sum_{k=k_{0}}^{\infty}\left|{a_{k}}\right|$. Obviously such a sequence is equivalent to the $\ell_{1}$-basis.
Corollary 5.9.
Let $M$ be a complete pointed metric space. For every $n\in\mathbb{N}$, let $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\in S_{\mathcal{F}(M)}$ such that
$\sum_{k}d(x_{k}^{n},y_{k}^{n})\underset{n\to\infty}{\longrightarrow}0\quad\text{and}\quad\sup_{n}\sum_{k}|\lambda_{k}^{n}|<\infty.$
Then there is an infinite $\mathbb{M}\subset\mathbb{N}$ such that $(\gamma_{n})_{n\in\mathbb{M}}$ is an asymptotically isometric $\ell_{1}$-basis.
Proof.
Let $(\varepsilon_{n})$ be a sequence of positive numbers such that $\prod_{n=1}^{\infty}(1-\varepsilon_{n})>0$. A routine argument employing the conclusion of Theorem 5.8 allows us to choose $(\gamma_{n_{k}})$ iteratively so that for every $x\in\operatorname{span}\left\{{\gamma_{n_{1}},\ldots,\gamma_{n_{k}}}\right\}$ one has $\left\|{x+a\gamma_{n_{k+1}}}\right\|\geqslant(1-\varepsilon_{n})(\left\|{x}\right\|+\left|{a}\right|)$. It follows that for every $k_{0}$ one has
$\left\|{\sum_{k=k_{0}}^{\infty}a_{k}\gamma_{n_{k}}}\right\|\geqslant\prod_{n=k_{0}}^{\infty}(1-\varepsilon_{n})\sum_{k=k_{0}}^{\infty}\left|{a_{k}}\right|.$
Since $\prod_{n=k_{0}}^{\infty}(1-\varepsilon_{n})\to 1$ as $k_{0}\to\infty$ we get the desired conclusion. ∎
Remark 5.10.
  1. (1)
    Even if $(\gamma_{n})$ is as in Corollary 5.9, the set $\left\{{\gamma_{n}:n\in\mathbb{N}}\right\}$ can still be uniformly regular. Indeed, denote by $(x_{i}^{n},y_{i}^{n})\subset[0,1]$, $i=1,\ldots,2^{n-1}$ the intervals removed in the $n^{\text{th}}$ step of the construction of the middle-thirds Cantor set. Now, consider the sequence $(\gamma_{n})\subset\mathcal{F}(M)$ defined as
    $\gamma_{n}:=\frac{1}{2^{n-1}}\sum_{i=1}^{2^{n-1}}m_{x_{i}^{n}y_{i}^{n}}.$
    Then following the proof of Proposition 3.3 one can show that $\left\{{\gamma_{n}:n\in\mathbb{N}}\right\}$ is uniformly regular.
  2. (2)
    We do not know, on the other hand, whether $(\gamma_{n})$ being as in Corollary 5.9 implies that $\left\{{\gamma_{n}:n\in\mathbb{N}}\right\}$ cannot be a V*-set. It seems reasonable to look for a counterexample in the space $\mathcal{F}(C[0,1])$ as $C[0,1]$ has no complemented copy of $\ell_{1}$ in it.
  3. (3)
    Given $n\in\mathbb{N}$, let $(x_{k}^{n})\subset[0,1]$, for $k=0,\ldots,2^{n}$, be the $n$-th generation dyadic numbers, and let $y_{k}^{n}=x_{k}^{n}+2^{-(n+k)}$. Then $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ with $n\to\infty$ but for any infinite $\mathbb{M}\subset\mathbb{N}$ we have $0,1]\subset\overline{\bigcup_{\mathbb{M}}\operatorname{supp}\gamma_{n}}$. (Here the exact definition of $\gamma_{n}$ is inconsequential, it can be for example $\gamma_{n}=\frac{1}{2^{n}+1}\sum_{k=0}^{2^{n}}m_{x_{k}^{n}y_{k}^{n}}$.) This shows that there are sequences $(\gamma_{n})$ as in Corollary 5.9 without any subsequences living in the free space of a complete purely 1-unrectifiable metric space.
The following corollary which provides another necessary condition for weak precompactness is strongly related to condition (v) of Proposition 4.6.
Corollary 5.11.
Let $M$ be a complete pointed metric space and let $W\subset\mathcal{F}({M})$ be weakly precompact. Then for every $\varepsilon>0$, there exists $\delta>0$ such that
$\operatorname{dist}\left(\gamma,\mathrm{CSM}_{\delta}\left({M}\right)\right)\geq\frac{\left\|{\gamma}\right\|}{2}-\varepsilon$
for every $\gamma\in W$.
For the proof we will need the following “big-perturbation lemma” for $\ell_{1}$-bases.
Lemma 5.12.
Let $(e_{n})\subset S_{X}$ be $C$-equivalent to the $\ell_{1}$ basis, i.e. $C\sum|a_{i}|\leqslant\|\sum a_{i}e_{i}\|$. Let $\|y_{n}-e_{n}\|\leqslant D<C$. Then $(y_{n})$ is equivalent to the $\ell_{1}$ basis.
Proof.
We have $\|\sum a_{n}y_{n}\|\geqslant\|\sum a_{n}e_{n}\|-\|\sum a_{n}(y_{n}-e_{n})\|\geqslant(C-D)\sum|a_{n}|$ and $\|\sum a_{n}y_{n}\|\leqslant\sum|a_{n}|\|y_{n}\|\leqslant(1+D)\sum|a_{n}|$. ∎
Proof of Corollary 5.11.
We prove the contrapositive. Assume that there exists $\varepsilon>0$ such that for every $n\in\mathbb{N}$, there exists $\gamma_{n}\in W$ such that
$\operatorname{dist}\left(\gamma_{n},\mathrm{CSM}_{\frac{1}{n}}\left({M}\right)\right)<\frac{\left\|{\gamma_{n}}\right\|}{2}-\varepsilon.$
We assume that $W$ is bounded, otherwise it is not weakly precompact and we are done. Let $R=\sup_{\gamma\in W}\left\|{\gamma}\right\|$. Let $\mu_{n}\in\mathrm{CSM}_{\frac{1}{n}}\left({M}\right)$ be such that $\left\|{\gamma_{n}-\mu_{n}}\right\|<\frac{\left\|{\gamma_{n}}\right\|}{2}-\varepsilon$. We show that $(\gamma_{n})$ has a subsequence equivalent to the $\ell_{1}$ basis and thus $W$ is not weakly precompact. First observe that $2\varepsilon<\left\|{\gamma_{n}}\right\|\leqslant R$, so we may instead find a subsequence of $(\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|})$ equivalent to the $\ell_{1}$-basis. Since $2\varepsilon<\left\|{\mu_{n}}\right\|$ for every $n$, Corollary 5.9 yields a subsequence of $(\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|})$ which is an asymptotically isometric $\ell_{1}$-basis. Thus if we find $D<1$ such that $\left\|{\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|}-\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|}}\right\|\leqslant D$ for every $n$, Lemma 5.12 will lead to the desired conclusion. But we have by the classical Massera-Schäffer inequality [19, Tapa 15]
$\left\|{\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|}-\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|}}\right\|\leqslant\frac{2\left\|{\gamma_{n}-\mu_{n}}\right\|}{\max\left\{{\left\|{\gamma_{n}}\right\|,\left\|{\mu_{n}}\right\|}\right\}}\leqslant\frac{\left\|{\gamma_{n}-\mu_{n}}\right\|}{\left\|{\gamma_{n}-\mu_{n}}\right\|+\varepsilon}\leqslant 1-\frac{2\varepsilon}{R}$
so setting $D=1-\frac{2\varepsilon}{R}$ finishes the proof. ∎
6. Final remarks
[…]
Proposition 6.1.
Let $M$ be a complete metric space and $K\subset\mathcal{F}(M)$ be bounded. Then the following assertions are equivalent:
  1. $(i)$
    $K$ is relatively compact.
  2. $(ii)$
[…]
Proof.
$(i)\implies(ii)$: Fix $\varepsilon>0$ and an open cover $K\subset\bigcup_{i=1}^{n}B(m_{i},\varepsilon)$, with $m_{i}$ being finitely supported for every $i\in\{1,\ldots,n\}$. Each $m_{i}$ can be written as a convex sum of molecules $m_{i}=\sum_{k=1}^{n_{i}}a_{(x_{k}^{i},y_{k}^{i})}m_{x_{k}^{i}y_{k}^{i}}$. Let $\delta:=\underset{i,k}{\min}\,d(x_{k}^{i},y_{k}^{i})$. Then assertion $(ii)$ clearly follows with $c=1$ and
$C=\{x_{k}^{i}:1\leqslant k\leqslant n_{i},\,1\leqslant i\leqslant n\}\cup\{y_{k}^{i}:1\leqslant k\leqslant n_{i},\,1\leqslant i\leqslant n\}.$
[…]
$a=\sum_{d(x,y)\geqslant\delta}a_{(x,y)}e_{(x,y)}+\sum_{d(x,y)<\delta}a_{(x,y)}e_{(x,y)}.$
If we denote $\widetilde{C}_{\delta}=\{(x,y)\in\widetilde{C}:d(x,y)\geqslant\delta\}$ then it is clear that
$Qa\in Q(c\left\|{\mu}\right\|\cdot B_{\ell_{1}(\widetilde{C}_{\delta})})+\varepsilon B_{\mathcal{F}(M)}.$
Hence $K\subset Q(cRB_{\ell_{1}(\widetilde{C}_{\delta})})+2\varepsilon B_{\mathcal{F}(M)}$ where $R=\max\left\{{\left\|{\mu}\right\|:\mu\in K}\right\}$. Now the fact that $Q(cRB_{\ell_{1}(\widetilde{C}_{\delta})})$ is relatively compact follows from the next lemma. ∎
Lemma 6.2.
Let $C\subset M$ be compact. For $\delta>0$ denote $\widetilde{C}_{\delta}=\{(x,y)\in\widetilde{C}:d(x,y)\geqslant\delta\}$. Then $Q\mathord{\upharpoonright}_{\ell_{1}(\widetilde{C}_{\delta})}$ is a compact operator.
Proof.
Since $Q(B_{\ell_{1}(\widetilde{C}_{\delta})})\,\subset\,\overline{\operatorname{conv}}\{Q(e_{x,y}):(x,y)\in\widetilde{C}_{\delta}\}$, it is enough to show that $\{Q(e_{(x,y)}):(x,y)\in\widetilde{C}_{\delta}\}$ is compact. Indeed, let $(x_{n},y_{n})\subset\widetilde{C}_{\delta}$. By compactness of $C$ and passing to a subsequence if necessary, we may assume that $(x_{n})$ and $(y_{n})$ converge to $x$ and $y$ respectively. Clearly $d(x,y)\geqslant\delta$ and so $Q(e_{x_{n},y_{n}})=m_{x_{n}y_{n}}\to m_{xy}=Q(e_{(x,y)})$. ∎
Remark 6.3.
We wish to highlight an alternative characterization of compact sets in Lipschitz free spaces, as established in [12]. To set the stage, recall Grothendieck’s compactness principle [17, p. 112], which states that every norm-compact subset of a Banach space is contained in the closed convex hull of a norm-null sequence. Adapting this principle to the setting of Lipschitz free spaces requires additional work, as demonstrated in [12, Corollary 3.5] (albeit in a slightly different context). Specifically, a closed subset $K\subset\mathcal{F}(M)$ is compact if and only if there exist sequences $(\alpha_{n})_{n}\in c_{0}$ (with $\alpha_{n}>0$) and $\big((x_{n},y_{n})\big)_{n}\subset\widetilde{M}$ such that
$K\subset\overline{\mathrm{conv}}\{\alpha_{n}m_{x_{n}y_{n}}:n\in\mathbb{N}\}.$
This characterization once again highlights the pivotal role of elementary molecules.
Acknowledgments
This work was initiated while the first two named authors were invited by Professor E. Pernecká at the Czech Technical University in Prague in 2023. They are grateful to Professor E. Pernecká for valuable discussions and for providing excellent working conditions during their visit. The authors also wish to thank Professor A. Quilis for useful conversations. The first and last named authors visited Université de Franche-Comté/Université Marie et Louis Pasteur on several occasions, for which they wish to express their gratitude.
R. J. Aliaga was partially supported by Grant PID2021-122126NB-C33 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU.
This work was partially supported by the French ANR projects ANR-20-CE40-0006 and ANR-24-CE40-0892-01.
T. Veeorg was supported by the Estonian Research Council grant (PRG1901).
A. Procházka and T. Veeorg were partially supported by the PARROT French-Estonian science and technology cooperation programme.